### v1

# The incompatibility of the Condorcet winner and loser criteria with positive involvement and resolvability

Abstract
We prove that there is no preferential voting method satisfying the Condorcet winner and loser criteria, positive involvement (if a candidate $x$ wins in an initial preference profile, then adding a voter who ranks $x$ uniquely first cannot cause $x$ to lose), and resolvability (if $x$ initially ties for winning, then $x$ can be made the unique winner by adding a single voter). In a previous note, we proved an analogous result assuming an additional axiom of ordinal margin invariance, which we now show is unnecessary for an impossibility theorem, at least if the desired voting method is defined for five-candidate elections.
1 Introduction
In the theory of preferential voting, where voters submit rankings of the candidates, there is a fundamental tradeoff when selecting a voting method between ensuring desired variable-voter properties, ensuring desired variable-candidate properties, and ensuring that ties are not too frequent. In this note, we prove an impossibility theorem illustrating this tradeoff. Here the relevant variable-voter axiom is that of positive involvement (see Saari 1995, Pérez 2001, Holliday and Pacuit 2021): if a candidate $x$ wins in an initial preference profile, then adding a voter who ranks $x$ uniquely first cannot cause $x$ to lose (see Definition 1.3). According to Pérez (2001, p. 605), this “may be seen as the minimum to require concerning the coherence in the winning set when new voters are added.” [^1] This seems too strong, however, as a more plausible minimum would be bullet-vote positive involvement: adding a voter who ranks a would-be winner $x$ uniquely first—and leaves all other candidates unranked or ranks them as tied after xx—cannot cause $x$ to lose. Another weaker condition is positive-negative involvement: adding a voter who ranks a winner $x$ uniquely first and a loser $y$ uniquely last cannot cause $x$ to lose and $y$ to win. On the candidate side, although the famous Condorcet winner and loser criteria (Definition 1.1-2) are not explicitly variable-candidate axioms, they can be motivated by considering candidates dropping out or joining the race: one might think that if $x$ would win (resp. lose) in every two-candidate sub-election, then $x$ should still win (resp. lose) in the full election—or else the election outcome might be seen as too contingent on which other losing candidates happened to stay in the race. [^2] As in the case of positive involvement, so too in the case of the Condorcet criteria—especially the Condorcet loser criterion—the motivation mentioned in the main text is not meant to be decisive. Finally, for a formalization of the idea that a voting method should not make ties too frequent, here we use the resolvability axiom that if $x$ is initially tied for winning, then $x$ can be made the unique winner by adding a single voter (Definition 1.4). As Tideman (1986, p. 25) puts it, “Resolvability may be thought of as a condition that ties happen only “by coincidence”, every tie being within one vote of being broken.”
2 Setup
Given infinite sets $\mathcal{V}$ and $\mathcal{X}$ of possible voters and possible candidates, respectively, a profile $\mathbf{P}$ is a function assigning to each voter $i$ in some nonempty finite set $V(\mathbf{P})\subseteq\mathcal{V}$ a strict weak order $\mathbf{P}(i)$ on some nonempty finite set $X(\mathbf{P})\subseteq\mathcal{X}$ of candidates. We call $\mathbf{P}$ linear if for each $i\in V(\mathbf{P})$, $\mathbf{P}(i)$ is a linear order, so there are no ties in $i$’s ranking. All results to follow continue to hold if all profiles are linear. If $\mathbf{P}$ and $\mathbf{Q}$ are profiles with $V(\mathbf{P})\cap V(\mathbf{Q})=\varnothing$ and $X(\mathbf{P})=X(\mathbf{Q})$, then $\mathbf{P}+\mathbf{Q}$ is the profile with $V(\mathbf{P}+\mathbf{Q})=V(\mathbf{P})\cup V(\mathbf{Q})$ defined in the obvious way: if $i\in V(\mathbf{P})$, then $(\mathbf{P}+\mathbf{Q})(i)=\mathbf{P}(i)$, and if $i\in V(\mathbf{Q})$, then $(\mathbf{P}+\mathbf{Q})(i)=\mathbf{Q}(i)$.
Given $x,y\in X(\mathbf{P})$, the margin of $x$ vs. $y$ in $\mathbf{P}$ is the number of voters who rank $x$ above $y$ minus the number who rank $y$ above $x$: $\text{Margin}_{\mathbf{P}}(x,y)=|\{i\in V(\mathbf{P})\mid(x,y)\in\mathbf{P}(i)\}|-|\{i\in V(\mathbf{P})\mid(y,x)\in\mathbf{P}(i)\}|$. The margin graph $\mathcal{M}(\mathbf{P})$ of $\mathbf{P}$ is the weighted directed graph whose set of vertices is $X(\mathbf{P})$ with an edge from $x$ to $y$ when $\text{Margin}_{\mathbf{P}}(x,y)>0$, weighted by $\text{Margin}_{\mathbf{P}}(x,y)$.
A voting method is a function $F$ assigning to each profile $\mathbf{P}$ a nonempty subset $F(\mathbf{P})\subseteq X(\mathbf{P})$. Note that this builds an assumption of “universal domain” into the definition of a voting method.
Definition 1.
Let $F$ be a voting method.
  1. 1.
    $F$ satisfies the Condorcet winner criterion (Condorcet 1785) if for every profile $\mathbf{P}$ and $x\in X(\mathbf{P})$, if $\text{Margin}_{\mathbf{P}}(x,y)>0$ for all $y\in X(\mathbf{P})\setminus\{x\}$ (so $x$ is a Condorcet winner), then $F(\mathbf{P})=\{x\}$.
  2. 2.
    $F$ satisfies the Condorcet loser criterion (Condorcet 1785) if for every profile $\mathbf{P}$ and $x\in X(\mathbf{P})$, if $\text{Margin}_{\mathbf{P}}(y,x)>0$ for all $y\in X(\mathbf{P})\setminus\{x\}$ (so $x$ is a Condorcet loser), then $x\not\in F(\mathbf{P})$.
  3. 3.
    $F$ satisfies positive involvement (Saari 1995) if for every profile $\mathbf{P}$, $x\in X(\mathbf{P})$, and profile $\mathbf{P}^{x}$ with $V(\mathbf{P})\cap V(\mathbf{P}^{x})=\varnothing$ and $X(\mathbf{P})=X(\mathbf{P}^{x})$, if $x\in F(\mathbf{P})$ and $\mathbf{P}^{x}$ consists of a single voter who ranks $x$ uniquely first, then $x\in F(\mathbf{P}+\mathbf{P}^{x})$.
  4. 4.
    $F$ satisfies resolvability (Tideman 1986) if for every profile $\mathbf{P}$, if $|F(\mathbf{P})|>1$, then for each $x\in F(\mathbf{P})$, there is some single-voter profile $\mathbf{S}$ with $V(\mathbf{P})\cap V(\mathbf{S})=\varnothing$ and $X(\mathbf{P})=X(\mathbf{S})$ such that $F(\mathbf{P}+\mathbf{S})=\{x\}$. [^3] In Holliday 2024, we used a weaker version of this axiom in the impossibility theorem, since there the weaker version sufficed: if $|F(\mathbf{P})|>1$, then there is some single-voter profile $\mathbf{S}$ such $|F(\mathbf{P}+\mathbf{S})|=1$. But Tideman’s original axiom (here stated in a variable-voter setting) requires that if $|F(\mathbf{P})|>1$, then for each $x\in F(\mathbf{P})$, there is some single-voter profile $\mathbf{S}$ such that $F(\mathbf{P}+\mathbf{S})=\{x\}$. This version facilitates the proof in this paper. In Holliday 2024, we also discussed asymptotic resolvability, but without the assumption of ordinal margin invariance made in that paper, asymptotic resolvability is more difficult to work with, so we do not discuss it here.
Each of the axioms in Definition 1 is independent of the others. The Borda count (Borda 1781) satisfies all of the axioms except the Condorcet winner criterion. Minimax (Simpson 1969, Kramer 1977) satisfies all of them except the Condorcet loser criterion. [^4] Minimax does not satisfy resolvability if we restrict its domain to linear profiles (so the added voter must submit a linear order), but the related leximax method (see Brandt et al. 2025), wherein the winning candidates are those whose vectors of margins, sorted in increasing order, are lexicographically largest, does satisfy resolvability when its domain is restricted (or not restricted) to linear profiles, as well as the Condorcet winner criterion and positive involvement (for all profiles). Ranked Pairs (Tideman 1987) satisfies all of them except positive involvement. Split Cycle (Holliday and Pacuit 2023) satisfies all of them except resolvability.
Below we will prove that there is no voting method satisfying all of the axioms in Definition 1.555Among the corollaries of our main theorem is that there is no voting method satisfying the Condorcet winner and loser criteria, negative involvement (Fishburn and Brams 1983, Pérez 2001) (if $x\not\in F(\mathbf{P})$ and $\mathbf{P}^{\prime}$ consists of a single voter who ranks $x$ uniquely last, then $x\not\in F(\mathbf{P}+\mathbf{P}^{\prime})$), resolvability, and neutral reversal (Saari 2003) (adding two voters with reversed linear orders does not change the set of winners); this uses the fact that if a voting method satisfies neutral reversal, then it satisfies positive involvement if and only if it satisfies negative involvement (see Ding et al. 2025, Prop. 3.18). The proof involves reasoning about five-candidate profiles. If we restrict the domain of voting methods to two or three-candidate profiles, then the axioms are jointly satisfied by the Minimax (Simpson 1969, Kramer 1977) voting method. For four candidates, the axioms are jointly inconsistent with an additional axiom of ordinal margin invariance (Holliday 2024), but we leave open whether they are jointly satisfiable without ordinal margin invariance if we only ask for a voting method defined for four-candidate profiles.
The impossibility theorem to follow makes use of the following concept and lemma.
Definition 2.
Given a profile $\mathbf{P}$, the defensible set $D(\mathbf{P})$ of $\mathbf{P}$ is the set of candidates $x\in X(\mathbf{P})$ such that for all $y\in X(\mathbf{P})$, there is a $z\in X(\mathbf{P})$ such that $\text{Margin}_{\mathbf{P}}(z,y)\geq\text{Margin}_{\mathbf{P}}(y,x)$.
The following lemma is essentially Lemma 3 of Pérez 1995 (cf. Lemma 1 of Pérez 2001), which is in turn an adaptation of a similar result of Moulin 1988, but we include a proof here for convenience.
Lemma 3 (Moulin 1988, Pérez 1995).
Let $F$ be a voting method satisfying positive involvement and the Condorcet winner criterion. Let $\mathbf{P}$ be a profile in which for all $x,x^{\prime},y,y^{\prime}\in X(\mathbf{P})$ with $x\neq x^{\prime}$ and $y\neq y^{\prime}$, if $\text{Margin}_{\mathbf{P}}(x,x^{\prime})>\text{Margin}_{\mathbf{P}}(y,y^{\prime})$, then $\text{Margin}_{\mathbf{P}}(x,x^{\prime})>\text{Margin}_{\mathbf{P}}(y,y^{\prime})+1$. Then $F(\mathbf{P})\subseteq D(\mathbf{P})$.
Proof.
Let $F$ be such a method, and suppose $x\in F(\mathbf{P})$. For contradiction, suppose $x\not\in D(\mathbf{P})$. Then there is a $y\in X(\mathbf{P})$ such that for every $z\in X(\mathbf{P})$, $\text{Margin}_{\mathbf{P}}(z,y)<\text{Margin}_{\mathbf{P}}(y,x)$, which implies $y\neq x$. Let $k=\max\{\text{Margin}_{\mathbf{P}}(z,y)\mid z\in X(\mathbf{P}),z\neq y\}$, so $k<\text{Margin}_{\mathbf{P}}(y,x)$. If $k<0$, then $y$ is the Condorcet winner in $\mathbf{P}$, so $F(\mathbf{P})=\{y\}$ by the Condorcet winner criterion, which contradicts $x\in F(\mathbf{P})$. Thus, $k\geq 0$. By our assumption about $\mathbf{P}$ in the lemma, $k+1<\text{Margin}_{\mathbf{P}}(y,x)$. Let $\mathbf{P}^{\prime}$ be obtained from $\mathbf{P}$ by adding $k+1$ voters who rank $x$ uniquely first and $y$ uniquely second, followed by any linear order of $X(\mathbf{P})\setminus\{x,y\}$. Hence by repeated application of positive involvement, $x\in F(\mathbf{P}^{\prime})$. But $y$ is the Condorcet winner in $\mathbf{P}^{\prime}$, so $F(\mathbf{P}^{\prime})=\{y\}$ by the Condorcet winner criterion, which contradicts $x\in F(\mathbf{P}^{\prime})$. Thus, $x\in D(\mathbf{P})$.∎
3 Result
Theorem 4.
There is no voting method satisfying the Condorcet winner and loser criteria, positive involvement, and resolvability.
Proof.
Suppose for contradiction that there is such a voting method $F$. Let $\mathbf{P}_{1}$ be the profile in Figure 1, whose margin graph is $\mathcal{M}_{1}$ in Figure 2.666Rather than using this specific $\mathbf{P}_{1}$, one may reason more abstractly as follows. By Debord’s Theorem (Debord 1987), there is a profile $\mathbf{Q}$ whose margin graph is $\mathcal{M}_{1}$. Now let $\mathbf{Q}^{+}$ be obtained from $\mathbf{Q}$ by adding, for each ranking of the five candidates, $10$ new voters with that ranking; then the margin graph of $\mathbf{Q}^{+}$ is still $\mathcal{M}_{1}$, and $\mathbf{Q}^{+}$ contains all the rankings needed for the rest of the argument, so one may use $\mathbf{Q}^{+}$ in place of $\mathbf{P}_{1}$. Then let $\mathbf{P}_{2},\dots,\mathbf{P}_{5}$ be obtained from $\mathbf{P}_{1}$ by adding and subtracting the profiles shown in the equations in Figure 2: e.g., $\mathbf{P}_{2}$ is obtained from $\mathbf{P}_{1}$ by adding 28 new voters who rank $a$ above $d$, $d$ above $b$, etc.; $\mathbf{P}_{3}$ is obtained from $\mathbf{P}_{2}$ by subtracting $8$ voters who rank $d$ above $a$, $a$ above $e$, etc. Note there are indeed such $8$ voters in $\mathbf{P}_{1}$, and for the purposes of $\mathbf{P}_{5}$, there are $10$ voters in $\mathbf{P}_{1}$ who rank $d$ above $a$, $a$ above $c$, etc. Then $\mathcal{M}_{2},\dots,\mathcal{M}_{5}$ are the margin graphs of $\mathbf{P}_{2},\dots,\mathbf{P}_{5}$, respectively.
67 39 37 24 21 20 13 10 10 8 6 2
$a$ $b$ $e$ $b$ $c$ $e$ $e$ $d$ $d$ $d$ $b$ $e$
$c$ $a$ $b$ $a$ $e$ $d$ $d$ $a$ $c$ $a$ $d$ $b$
$d$ $c$ $d$ $c$ $b$ $b$ $c$ $c$ $e$ $e$ $a$ $d$
$e$ $e$ $c$ $d$ $d$ $a$ $b$ $b$ $b$ $c$ $c$ $a$
$b$ $d$ $a$ $e$ $a$ $c$ $a$ $e$ $a$ $b$ $e$ $c$
Figure 1: the profile $\mathbf{P}_{1}$ used in the proof of Theorem 4
$\mathcal{M}_{1}$
aabbccddee87879595113311454551519999979777
+\;+ 28 $a$ $d$ $b$ $e$ $c$ ==\;\; $\mathcal{M}_{2}$aabbccddee595912312327273131272717177979717169692121 == 8 $d$ $a$ $e$ $c$ $b$ ++\;\; $\mathcal{M}_{3}$aabbccddee676711511535353939191925257171636377771313
$\mathcal{M}_{3}$
aabbccddee676711511535353939191925257171636377771313
+\;+ 26 $b$ $d$ $e$ $c$ $a$ ==\;\; $\mathcal{M}_{4}$aabbccddee939389896161131377114545373751513939 == 10 $d$ $a$ $c$ $b$ $e$ ++\;\; $\mathcal{M}_{5}$aabbccddee1031037979717123231717993535474741412929
Figure 2: margin graphs used in the proof of Theorem 4 with defensible candidates shaded gray
Now in $\mathbf{P}_{1}$, $a$ is the only defensible candidate who is not a Condorcet loser (see $\mathcal{M}_{1}$ in Figure 2). Thus, by the Condorcet loser criterion and Lemma 3, $F(\mathbf{P}_{1})=\{a\}$. Since $\mathbf{P}_{2}$ is obtained from $\mathbf{P}_{1}$ by adding voters all of whom rank $a$ uniquely first, $F(\mathbf{P}_{1})=\{a\}$ and repeated application of positive involvement together imply $a\in F(\mathbf{P}_{2})$. Then by single-voter resolvability, there is some “empty” [^7] Here $\mathbf{S}_{2}$ is not literally an empty profile; if we already have $F(\mathbf{P}_{2})=\{a\}$, then let $\mathbf{P}_{2}+\mathbf{S}_{2}$’ denote $\mathbf{P}_{2}$, $\mathbf{P}_{3}+\mathbf{S}_{2}$’ denote $\mathbf{P}_{3}$, etc. The analogous convention applies to $\mathbf{S}_{4}$ below if $F(\mathbf{P}_{4}+\mathbf{S}_{2})=\{b\}$. or single-voter profile $\mathbf{S}_{2}$ such that $F(\mathbf{P}_{2}+\mathbf{S}_{2})=\{a\}$}. Since $\mathbf{P}_{2}+\mathbf{S}_{2}$ is obtained from $\mathbf{P}_{3}+\mathbf{S}_{2}$ by adding voters all of whom rank $d$ uniquely first, $F(\mathbf{P}_{2}+\mathbf{S}_{2})=\{a\}$} and positive involvement together imply $d\not\in F(\mathbf{P}_{3}+\mathbf{S}_{2})$. Since any difference between weights of two distinct edges in $\mathcal{M}_{3}$ is of size at least $4$, no matter which single-voter (or empty) profile $\mathbf{S}_{2}$ is, $b$ is the only defensible candidate besides $d$ in $\mathbf{P}_{3}+\mathbf{S}_{2}$, and $\mathbf{P}_{3}+\mathbf{S}_{2}$ satisfies the margin separation assumption of Lemma 3, so $d\not\in F(\mathbf{P}_{3}+\mathbf{S}_{2})$ implies $F(\mathbf{P}_{3}+\mathbf{S}_{2})=\{b\}$ by Lemma 3. Then since $\mathbf{P}_{4}+\mathbf{S}_{2}$ is obtained from $\mathbf{P}_{3}+\mathbf{S}_{2}$ by adding voters all of whom rank $b$ uniquely first, $F(\mathbf{P}_{3}+\mathbf{S}_{2})=\{b\}$ and positive involvement together imply $b\in F(\mathbf{P}_{4}+\mathbf{S}_{2})$. Then by single-voter resolvability, there is some “empty” or single-voter profile $\mathbf{S}_{4}$ such that $F(\mathbf{P}_{4}+\mathbf{S}_{2}+\mathbf{S}_{4})=\{b\}$. Since $\mathbf{P}_{4}+\mathbf{S}_{2}+\mathbf{S}_{4}$ is obtained from $\mathbf{P}_{5}+\mathbf{S}_{2}+\mathbf{S}_{4}$ by adding voters all of whom rank $d$ uniquely first, $F(\mathbf{P}_{4}+\mathbf{S}_{2}+\mathbf{S}_{4})=\{b\}$ and positive involvement together imply $d\not\in F(\mathbf{P}_{5}+\mathbf{S}_{2}+\mathbf{S}_{4})$. Finally, since any difference between weights of two distinct edges in $\mathcal{M}_{5}$ is of size at least $6$, no matter which single-voter (or empty) profiles $\mathbf{S}_{2}$ and $\mathbf{S}_{4}$ are, $d$ is the only defensible candidate in $\mathbf{P}_{5}+\mathbf{S}_{2}+\mathbf{S}_{4}$, and $\mathbf{P}_{5}+\mathbf{S}_{2}+\mathbf{S}_{4}$ satisfies the margin separation assumption of Lemma 3, so $F(\mathbf{P}_{5}+\mathbf{S}_{2}+\mathbf{S}_{4})=\{d\}$ by Lemma 3. Thus, we have a contradiction.∎
Thus, we must learn to live with the non-existence of what we previously called the “holy grail” (Holliday 2024) of a voting method that satisfies the axioms in Theorem 4.
Acknowledgements
I thank Emma Boniface and Diego Bejarano for helpful discussions of the problem posed at the end of Holliday 2024.

### v2

@@
 Abstract
-We prove that there is no preferential voting method satisfying the Condorcet winner and loser criteria, positive involvement (if a candidate $x$ wins in an initial preference profile, then adding a voter who ranks $x$ uniquely first cannot cause $x$ to lose), and resolvability (if $x$ initially ties for winning, then $x$ can be made the unique winner by adding a single voter).
+We prove that there is no preferential voting method satisfying the Condorcet winner and loser criteria, positive involvement (if a candidate $x$ wins in an initial preference profile, then adding a voter who ranks $x$ uniquely first cannot cause $x$ to lose), and $n$-voter resolvability (if $x$ initially ties for winning, then $x$ can be made the unique winner by adding some set of up to $n$ voters).
+This impossibility theorem holds for any positive integer $n$.
 In a previous note, we proved an analogous result assuming an additional axiom of ordinal margin invariance, which we now show is unnecessary for an impossibility theorem, at least if the desired voting method is defined for five-candidate elections.
@@
 [^2] As in the case of positive involvement, so too in the case of the Condorcet criteria—especially the Condorcet loser criterion—the motivation mentioned in the main text is not meant to be decisive.
-Finally, for a formalization of the idea that a voting method should not make ties too frequent, here we use the resolvability axiom that if $x$ is initially tied for winning, then $x$ can be made the unique winner by adding a single voter (Definition 1.4).
+Finally, for a formalization of the idea that a voting method should not make ties too frequent, here we use the resolvability axiom that if $x$ is initially tied for winning, then $x$ can be made the unique winner by adding a new voter (Definition 1.4).
 As Tideman (1986, p.
-25) puts it, “Resolvability may be thought of as a condition that ties happen only “by coincidence”, every tie being within one vote of being broken.”
+25) puts it, “Resolvability may be thought of as a condition that ties happen only “by coincidence”, every tie being within one vote of being broken.” In fact, we will ultimately consider, for any positive integer $n$, the axiom of $n$-voter resolvability, whereby any tied winner can be made the unique winner by adding some set of up to $n$ new voters.
 2 Setup
@@
 For four candidates, the axioms are jointly inconsistent with an additional axiom of ordinal margin invariance (Holliday 2024), but we leave open whether they are jointly satisfiable without ordinal margin invariance if we only ask for a voting method defined for four-candidate profiles.
-The impossibility theorem to follow makes use of the following concept and lemma.
+The impossibility proof to follow makes use of the following concept and lemma.
 Definition 2.
@@
 Then $\mathcal{M}_{2},\dots,\mathcal{M}_{5}$ are the margin graphs of $\mathbf{P}_{2},\dots,\mathbf{P}_{5}$, respectively.
-67 39 37 24 21 20 13 10 10 8 6 2
-$a$ $b$ $e$ $b$ $c$ $e$ $e$ $d$ $d$ $d$ $b$ $e$
-$c$ $a$ $b$ $a$ $e$ $d$ $d$ $a$ $c$ $a$ $d$ $b$
-$d$ $c$ $d$ $c$ $b$ $b$ $c$ $c$ $e$ $e$ $a$ $d$
-$e$ $e$ $c$ $d$ $d$ $a$ $b$ $b$ $b$ $c$ $c$ $a$
-$b$ $d$ $a$ $e$ $a$ $c$ $a$ $e$ $a$ $b$ $e$ $c$
+63 39 35 24 21 22 11 10 14 8 6 4
+$a$ $b$ $e$ $b$ $c$ $e$ $e$ $d$ $d$ $d$ $b$ $a$
+$c$ $a$ $b$ $a$ $e$ $d$ $d$ $a$ $c$ $a$ $d$ $e$
+$d$ $c$ $d$ $c$ $b$ $b$ $c$ $c$ $e$ $e$ $a$ $c$
+$e$ $e$ $c$ $d$ $d$ $a$ $b$ $b$ $b$ $c$ $c$ $b$
+$b$ $d$ $a$ $e$ $a$ $c$ $a$ $e$ $a$ $b$ $e$ $d$
 Figure 1: the profile $\mathbf{P}_{1}$ used in the proof of Theorem 4
 $\mathcal{M}_{1}$
-aabbccddee87879595113311454551519999979777
-+\;+ 28 $a$ $d$ $b$ $e$ $c$ ==\;\; $\mathcal{M}_{2}$aabbccddee595912312327273131272717177979717169692121 == 8 $d$ $a$ $e$ $c$ $b$ ++\;\; $\mathcal{M}_{3}$aabbccddee676711511535353939191925257171636377771313
+aabbccddee87879595553311454551519999979777
++\;+ 28 $a$ $d$ $b$ $e$ $c$ ==\;\; $\mathcal{M}_{2}$aabbccddee595912312323233131272717177979717169692121 == 8 $d$ $a$ $e$ $c$ $b$ ++\;\; $\mathcal{M}_{3}$aabbccddee676711511531313939191925257171636377771313
 $\mathcal{M}_{3}$
-aabbccddee676711511535353939191925257171636377771313
-+\;+ 26 $b$ $d$ $e$ $c$ $a$ ==\;\; $\mathcal{M}_{4}$aabbccddee939389896161131377114545373751513939 == 10 $d$ $a$ $c$ $b$ $e$ ++\;\; $\mathcal{M}_{5}$aabbccddee1031037979717123231717993535474741412929
+aabbccddee676711511531313939191925257171636377771313
++\;+ 26 $b$ $d$ $e$ $c$ $a$ ==\;\; $\mathcal{M}_{4}$aabbccddee939389895757131377114545373751513939 == 10 $d$ $a$ $c$ $b$ $e$ ++\;\; $\mathcal{M}_{5}$aabbccddee1031037979676723231717993535474741412929
 Figure 2: margin graphs used in the proof of Theorem 4 with defensible candidates shaded gray
@@
 Since $\mathbf{P}_{2}$ is obtained from $\mathbf{P}_{1}$ by adding voters all of whom rank $a$ uniquely first, $F(\mathbf{P}_{1})=\{a\}$ and repeated application of positive involvement together imply $a\in F(\mathbf{P}_{2})$.
-Then by single-voter resolvability, there is some “empty” [^7] Here $\mathbf{S}_{2}$ is not literally an empty profile; if we already have $F(\mathbf{P}_{2})=\{a\}$, then let $\mathbf{P}_{2}+\mathbf{S}_{2}$’ denote $\mathbf{P}_{2}$, $\mathbf{P}_{3}+\mathbf{S}_{2}$’ denote $\mathbf{P}_{3}$, etc.
+Then by resolvability, there is some “empty” [^7] Here $\mathbf{S}_{2}$ is not literally an empty profile; if we already have $F(\mathbf{P}_{2})=\{a\}$, then let $\mathbf{P}_{2}+\mathbf{S}_{2}$’ denote $\mathbf{P}_{2}$, $\mathbf{P}_{3}+\mathbf{S}_{2}$’ denote $\mathbf{P}_{3}$, etc.
 The analogous convention applies to $\mathbf{S}_{4}$ below if $F(\mathbf{P}_{4}+\mathbf{S}_{2})=\{b\}$.
@@
 Then since $\mathbf{P}_{4}+\mathbf{S}_{2}$ is obtained from $\mathbf{P}_{3}+\mathbf{S}_{2}$ by adding voters all of whom rank $b$ uniquely first, $F(\mathbf{P}_{3}+\mathbf{S}_{2})=\{b\}$ and positive involvement together imply $b\in F(\mathbf{P}_{4}+\mathbf{S}_{2})$.
-Then by single-voter resolvability, there is some “empty” or single-voter profile $\mathbf{S}_{4}$ such that $F(\mathbf{P}_{4}+\mathbf{S}_{2}+\mathbf{S}_{4})=\{b\}$.
+Then by resolvability, there is some “empty” or single-voter profile $\mathbf{S}_{4}$ such that $F(\mathbf{P}_{4}+\mathbf{S}_{2}+\mathbf{S}_{4})=\{b\}$.
 Since $\mathbf{P}_{4}+\mathbf{S}_{2}+\mathbf{S}_{4}$ is obtained from $\mathbf{P}_{5}+\mathbf{S}_{2}+\mathbf{S}_{4}$ by adding voters all of whom rank $d$ uniquely first, $F(\mathbf{P}_{4}+\mathbf{S}_{2}+\mathbf{S}_{4})=\{b\}$ and positive involvement together imply $d\not\in F(\mathbf{P}_{5}+\mathbf{S}_{2}+\mathbf{S}_{4})$.
 Finally, since any difference between weights of two distinct edges in $\mathcal{M}_{5}$ is of size at least $6$, no matter which single-voter (or empty) profiles $\mathbf{S}_{2}$ and $\mathbf{S}_{4}$ are, $d$ is the only defensible candidate in $\mathbf{P}_{5}+\mathbf{S}_{2}+\mathbf{S}_{4}$, and $\mathbf{P}_{5}+\mathbf{S}_{2}+\mathbf{S}_{4}$ satisfies the margin separation assumption of Lemma 3, so $F(\mathbf{P}_{5}+\mathbf{S}_{2}+\mathbf{S}_{4})=\{d\}$ by Lemma 3.
-Thus, we have a contradiction.∎
-Thus, we must learn to live with the non-existence of what we previously called the “holy grail” (Holliday 2024) of a voting method that satisfies the axioms in Theorem 4.
+Therefore, we have a contradiction.∎
+The same proof strategy can be used to show that for each positive integer $n$, resolvability can be replaced by $n$-voter resolvability: for every profile $\mathbf{P}$, if $|F(\mathbf{P})|>1$, then for each $x\in F(\mathbf{P})$, there is a profile $\mathbf{Q}$ with $|V(\mathbf{Q})|\leq n$, $V(\mathbf{P})\cap V(\mathbf{Q})=\varnothing$, and $X(\mathbf{P})=X(\mathbf{Q})$ such that $F(\mathbf{P}+\mathbf{Q})=\{x\}$.
+Theorem 5.
+For each positive integer $n$, there is no voting method satisfying the Condorcet winner and loser criteria, positive involvement, and $n$-voter resolvability.
+Proof.
+Replace each voter in the profiles $\mathbf{P}_{1},\dots,\mathbf{P}_{5}$ by $n$ copies of that voter, which results in a multiplication of all margins in $\mathcal{M}_{1},\dots,\mathcal{M}_{5}$ by $n$, and then the proof from Theorem 4 goes through as before but now with $|V(\mathbf{S}_{2})|\leq n$ and $|V(\mathbf{S}_{4})|\leq n$.∎
+Thus, for any method satisfying the Condorcet winner and loser criteria and positive involvement, no matter the size of $n$, sometimes a tied winner cannot be made a unique winner even by the addition of up to $n$ voters.
 Acknowledgements

### v3

@@
 1 Introduction
-In the theory of preferential voting, where voters submit rankings of the candidates, there is a fundamental tradeoff when selecting a voting method between ensuring desired variable-voter properties, ensuring desired variable-candidate properties, and ensuring that ties are not too frequent.
+In the theory of preferential voting, where voters submit rankings of the candidates, there is a fundamental tradeoff when selecting a voting method between ensuring desired variable-voter properties, ensuring desired variable-candidate properties, and ensuring that ties are not too common.
 In this note, we prove an impossibility theorem illustrating this tradeoff.
@@
 According to Pérez (2001, p.
-605), this “may be seen as the minimum to require concerning the coherence in the winning set when new voters are added.” [^1] This seems too strong, however, as a more plausible minimum would be bullet-vote positive involvement: adding a voter who ranks a would-be winner $x$ uniquely first—and leaves all other candidates unranked or ranks them as tied after xx—cannot cause $x$ to lose.
-Another weaker condition is positive-negative involvement: adding a voter who ranks a winner $x$ uniquely first and a loser $y$ uniquely last cannot cause $x$ to lose and $y$ to win.
+605), this “may be seen as the minimum to require concerning the coherence in the winning set when new voters are added.” [^1] This seems too strong, however, as a more plausible minimum would be bullet-vote positive involvement: adding a voter who ranks an initial winner $x$ uniquely first—and leaves all other candidates unranked or ranks them as tied after xx—cannot cause $x$ to lose.
+Another weaker condition is positive-negative involvement: adding a voter who ranks an initial winner $x$ uniquely first and an initial loser $y$ uniquely last cannot cause $x$ to lose and $y$ to win.
 On the candidate side, although the famous Condorcet winner and loser criteria (Definition 1.1-2) are not explicitly variable-candidate axioms, they can be motivated by considering candidates dropping out or joining the race: one might think that if $x$ would win (resp.
@@
 [^2] As in the case of positive involvement, so too in the case of the Condorcet criteria—especially the Condorcet loser criterion—the motivation mentioned in the main text is not meant to be decisive.
-Finally, for a formalization of the idea that a voting method should not make ties too frequent, here we use the resolvability axiom that if $x$ is initially tied for winning, then $x$ can be made the unique winner by adding a new voter (Definition 1.4).
+Finally, for a formalization of the idea that a voting method should not make ties too common, here we use the resolvability axiom that if $x$ is initially tied for winning, then $x$ can be made the unique winner by adding a single new voter (Definition 1.4).
 As Tideman (1986, p.
@@
 For four candidates, the axioms are jointly inconsistent with an additional axiom of ordinal margin invariance (Holliday 2024), but we leave open whether they are jointly satisfiable without ordinal margin invariance if we only ask for a voting method defined for four-candidate profiles.
-The impossibility proof to follow makes use of the following concept and lemma.
+The impossibility proof in the next section makes use of the following concept.
 Definition 2.
@@
 Suppose for contradiction that there is such a voting method $F$.
-Let $\mathbf{P}_{1}$ be the profile in Figure 1, whose margin graph is $\mathcal{M}_{1}$ in Figure 2.666Rather than using this specific $\mathbf{P}_{1}$, one may reason more abstractly as follows.
+Let $\mathbf{P}_{1}$ be a profile as in Figure 1, whose margin graph is $\mathcal{M}_{1}$ in Figure 2.666Rather than using this specific $\mathbf{P}_{1}$, one may reason more abstractly as follows.
 By Debord’s Theorem (Debord 1987), there is a profile $\mathbf{Q}$ whose margin graph is $\mathcal{M}_{1}$.
-Now let $\mathbf{Q}^{+}$ be obtained from $\mathbf{Q}$ by adding, for each ranking of the five candidates, $10$ new voters with that ranking; then the margin graph of $\mathbf{Q}^{+}$ is still $\mathcal{M}_{1}$, and $\mathbf{Q}^{+}$ contains all the rankings needed for the rest of the argument, so one may use $\mathbf{Q}^{+}$ in place of $\mathbf{P}_{1}$.
-Then let $\mathbf{P}_{2},\dots,\mathbf{P}_{5}$ be obtained from $\mathbf{P}_{1}$ by adding and subtracting the profiles shown in the equations in Figure 2: e.g., $\mathbf{P}_{2}$ is obtained from $\mathbf{P}_{1}$ by adding 28 new voters who rank $a$ above $d$, $d$ above $b$, etc.; $\mathbf{P}_{3}$ is obtained from $\mathbf{P}_{2}$ by subtracting $8$ voters who rank $d$ above $a$, $a$ above $e$, etc.
-Note there are indeed such $8$ voters in $\mathbf{P}_{1}$, and for the purposes of $\mathbf{P}_{5}$, there are $10$ voters in $\mathbf{P}_{1}$ who rank $d$ above $a$, $a$ above $c$, etc.
+Now let $\mathbf{Q}^{+}$ be obtained from $\mathbf{Q}$ by adding, for each ranking of the five candidates, $7$ new voters with that ranking; then the margin graph of $\mathbf{Q}^{+}$ is still $\mathcal{M}_{1}$, and $\mathbf{Q}^{+}$ contains all the rankings needed for the rest of the argument, so one may use $\mathbf{Q}^{+}$ in place of $\mathbf{P}_{1}$.
+Then let $\mathbf{P}_{2},\dots,\mathbf{P}_{5}$ be obtained from $\mathbf{P}_{1}$ by adding and subtracting the profiles shown in the equations in Figure 2: e.g., $\mathbf{P}_{2}$ is obtained from $\mathbf{P}_{1}$ by adding 26 new voters who rank $a$ above $d$, $d$ above $b$, etc.; $\mathbf{P}_{3}$ is obtained from $\mathbf{P}_{2}$ by subtracting $7$ voters who rank $d$ above $a$, $a$ above $c$, etc.
+Note there are indeed such $7$ voters in $\mathbf{P}_{1}$, and for the purposes of $\mathbf{P}_{5}$, there are $7$ voters in $\mathbf{P}_{1}$ who rank $d$ above $b$, $b$ above $a$, etc.
 Then $\mathcal{M}_{2},\dots,\mathcal{M}_{5}$ are the margin graphs of $\mathbf{P}_{2},\dots,\mathbf{P}_{5}$, respectively.
-63 39 35 24 21 22 11 10 14 8 6 4
-$a$ $b$ $e$ $b$ $c$ $e$ $e$ $d$ $d$ $d$ $b$ $a$
-$c$ $a$ $b$ $a$ $e$ $d$ $d$ $a$ $c$ $a$ $d$ $e$
-$d$ $c$ $d$ $c$ $b$ $b$ $c$ $c$ $e$ $e$ $a$ $c$
-$e$ $e$ $c$ $d$ $d$ $a$ $b$ $b$ $b$ $c$ $c$ $b$
-$b$ $d$ $a$ $e$ $a$ $c$ $a$ $e$ $a$ $b$ $e$ $d$
-Figure 1: the profile $\mathbf{P}_{1}$ used in the proof of Theorem 4
+69 64 46 20 18 2
+$d$ $e$ $b$ $c$ $d$ $e$
+$a$ $b$ $c$ $d$ $b$ $d$
+$c$ $a$ $a$ $e$ $a$ $c$
+$e$ $c$ $e$ $b$ $c$ $b$
+$b$ $d$ $d$ $a$ $e$ $a$
+Figure 1: the profile $\mathbf{P}_{1}$ used in the proof of Theorem 4. The number at the top of each column indicates the number of voters submitting the ranking in that column.
 $\mathcal{M}_{1}$
-aabbccddee87879595553311454551519999979777
-+\;+ 28 $a$ $d$ $b$ $e$ $c$ ==\;\; $\mathcal{M}_{2}$aabbccddee595912312323233131272717177979717169692121 == 8 $d$ $a$ $e$ $c$ $b$ ++\;\; $\mathcal{M}_{3}$aabbccddee676711511531313939191925257171636377771313
+aabbccddee8383114747818137371141418787919155
++\;+ 26 $a$ $d$ $b$ $e$ $c$ ==\;\; $\mathcal{M}_{2}$aabbccddee109109272773735555636315156161252521216565 == 7 $d$ $a$ $c$ $e$ $b$ ++\;\; $\mathcal{M}_{3}$aabbccddee102102343466666262707022225454181814145858
 $\mathcal{M}_{3}$
-aabbccddee676711511531313939191925257171636377771313
-+\;+ 26 $b$ $d$ $e$ $c$ $a$ ==\;\; $\mathcal{M}_{4}$aabbccddee939389895757131377114545373751513939 == 10 $d$ $a$ $c$ $b$ $e$ ++\;\; $\mathcal{M}_{5}$aabbccddee1031037979676723231717993535474741412929
-Figure 2: margin graphs used in the proof of Theorem 4 with defensible candidates shaded gray
+aabbccddee102102343466666262707022225454181814145858
++\;+ 23 $b$ $d$ $e$ $a$ $c$ ==\;\; $\mathcal{M}_{4}$aabbccddee12512511114343858593935531311137373535 == 7 $d$ $b$ $a$ $c$ $e$ ++\;\; $\mathcal{M}_{5}$aabbccddee1181181818363678788686121266242430304242
+Figure 2: margin graphs used in the proof of Theorem 4 with defensible candidates shaded gray.
+An edge from candidate $x$ to candidate $y$ labeled by $k$ indicates that the margin of $x$ vs.
+$y$ is $k$.
 Now in $\mathbf{P}_{1}$, $a$ is the only defensible candidate who is not a Condorcet loser (see $\mathcal{M}_{1}$ in Figure 2).
@@
 Therefore, we have a contradiction.∎
+Remark 5.
+Inspection of the proof of Theorem 4 shows that resolvability can be replaced by quasi-resoluteness (in the terminology of Holliday and Pacuit 2023): $|F(\mathbf{P})|=1$ for any profile $\mathbf{P}$ that is uniquely weighted in the sense that for all $x,y,x^{\prime},y^{\prime}\in X(\mathbf{P})$ with $x\neq y$ and $x^{\prime}\neq y^{\prime}$, if $(x,y)\neq(x^{\prime},y^{\prime})$, then $\text{Margin}_{\mathbf{P}}(x,y)\neq\text{Margin}_{\mathbf{P}}(x^{\prime},y^{\prime})$.
+It also shows that the package of positive involvement and resolvability can be replaced by strict positive involvement (in the terminology of Dominik Peters, personal communication), which is defined like positive involvement but with $F(\mathbf{P}+\mathbf{P}^{x})=\{x\}$ in place of $x\in F(\mathbf{P}+\mathbf{P}^{x})$.
 The same proof strategy can be used to show that for each positive integer $n$, resolvability can be replaced by $n$-voter resolvability: for every profile $\mathbf{P}$, if $|F(\mathbf{P})|>1$, then for each $x\in F(\mathbf{P})$, there is a profile $\mathbf{Q}$ with $|V(\mathbf{Q})|\leq n$, $V(\mathbf{P})\cap V(\mathbf{Q})=\varnothing$, and $X(\mathbf{P})=X(\mathbf{Q})$ such that $F(\mathbf{P}+\mathbf{Q})=\{x\}$.
@@
 Acknowledgements
-I thank Emma Boniface and Diego Bejarano for helpful discussions of the problem posed at the end of Holliday 2024.
+I thank Emma Boniface and Diego Bejarano for helpful discussions of the problem posed at the end of Holliday 2024 and Dominik Peters for helpful feedback on this paper.

### v4

@@
-The incompatibility of the Condorcet winner and loser criteria with positive involvement and resolvability
+The incompatibility of the Condorcet winner and loser criteria with positive or negative involvement and resolvability
 Abstract
@@
 This impossibility theorem holds for any positive integer $n$.
+In addition, positive involvement can be replaced by negative involvement (if a candidate $x$ loses in an initial preference profile, then adding a voter who ranks $x$ uniquely last cannot cause $x$ to win).
 In a previous note, we proved an analogous result assuming an additional axiom of ordinal margin invariance, which we now show is unnecessary for an impossibility theorem, at least if the desired voting method is defined for five-candidate elections.
@@
 Here the relevant variable-voter axiom is that of positive involvement (see Saari 1995, Pérez 2001, Holliday and Pacuit 2021): if a candidate $x$ wins in an initial preference profile, then adding a voter who ranks $x$ uniquely first cannot cause $x$ to lose (see Definition 1.3).
+Alternatively, we can use the axiom of negative involvement (see Fishburn and Brams 1983, Pérez 2001): if a candidate $x$ loses in an initial preference profile, then adding a voter who ranks $x$ uniquely last cannot cause $x$ to win.
 According to Pérez (2001, p.
-605), this “may be seen as the minimum to require concerning the coherence in the winning set when new voters are added.” [^1] This seems too strong, however, as a more plausible minimum would be bullet-vote positive involvement: adding a voter who ranks an initial winner $x$ uniquely first—and leaves all other candidates unranked or ranks them as tied after xx—cannot cause $x$ to lose.
-Another weaker condition is positive-negative involvement: adding a voter who ranks an initial winner $x$ uniquely first and an initial loser $y$ uniquely last cannot cause $x$ to lose and $y$ to win.
+605), these “may be seen as the minimum to require concerning the coherence in the winning set when new voters are added.” [^1] This seems too strong, however, as a more plausible minimum would be positive-negative involvement: adding a voter who ranks an initial winner $x$ uniquely first and an initial loser $y$ uniquely last cannot cause $x$ to lose and $y$ to win.
+Another weaker condition is bullet-vote positive involvement: adding a voter who ranks an initial winner $x$ uniquely first—and leaves all other candidates unranked or ranks them as tied after xx—cannot cause $x$ to lose.
 On the candidate side, although the famous Condorcet winner and loser criteria (Definition 1.1-2) are not explicitly variable-candidate axioms, they can be motivated by considering candidates dropping out or joining the race: one might think that if $x$ would win (resp.
@@
 Split Cycle (Holliday and Pacuit 2023) satisfies all of them except resolvability.
-Below we will prove that there is no voting method satisfying all of the axioms in Definition 1.555Among the corollaries of our main theorem is that there is no voting method satisfying the Condorcet winner and loser criteria, negative involvement (Fishburn and Brams 1983, Pérez 2001) (if $x\not\in F(\mathbf{P})$ and $\mathbf{P}^{\prime}$ consists of a single voter who ranks $x$ uniquely last, then $x\not\in F(\mathbf{P}+\mathbf{P}^{\prime})$), resolvability, and neutral reversal (Saari 2003) (adding two voters with reversed linear orders does not change the set of winners); this uses the fact that if a voting method satisfies neutral reversal, then it satisfies positive involvement if and only if it satisfies negative involvement (see Ding et al.
-2025, Prop.
-3.18).
+Below we will prove that there is no voting method satisfying all of the axioms in Definition 1.
 The proof involves reasoning about five-candidate profiles.
@@
 By Debord’s Theorem (Debord 1987), there is a profile $\mathbf{Q}$ whose margin graph is $\mathcal{M}_{1}$.
-Now let $\mathbf{Q}^{+}$ be obtained from $\mathbf{Q}$ by adding, for each ranking of the five candidates, $7$ new voters with that ranking; then the margin graph of $\mathbf{Q}^{+}$ is still $\mathcal{M}_{1}$, and $\mathbf{Q}^{+}$ contains all the rankings needed for the rest of the argument, so one may use $\mathbf{Q}^{+}$ in place of $\mathbf{P}_{1}$.
-Then let $\mathbf{P}_{2},\dots,\mathbf{P}_{5}$ be obtained from $\mathbf{P}_{1}$ by adding and subtracting the profiles shown in the equations in Figure 2: e.g., $\mathbf{P}_{2}$ is obtained from $\mathbf{P}_{1}$ by adding 26 new voters who rank $a$ above $d$, $d$ above $b$, etc.; $\mathbf{P}_{3}$ is obtained from $\mathbf{P}_{2}$ by subtracting $7$ voters who rank $d$ above $a$, $a$ above $c$, etc.
-Note there are indeed such $7$ voters in $\mathbf{P}_{1}$, and for the purposes of $\mathbf{P}_{5}$, there are $7$ voters in $\mathbf{P}_{1}$ who rank $d$ above $b$, $b$ above $a$, etc.
+Now let $\mathbf{Q}^{+}$ be obtained from $\mathbf{Q}$ by adding, for each ranking of the five candidates, 77 new voters with that ranking; then the margin graph of $\mathbf{Q}^{+}$ is still $\mathcal{M}_{1}$, and $\mathbf{Q}^{+}$ contains all the rankings needed for the rest of the argument, so one may use $\mathbf{Q}^{+}$ in place of $\mathbf{P}_{1}$.
+After the initial construction of this proof using another version of $\mathcal{M}_{1}$ and different numbers of rankings to add/subtract to obtain an appropriate sequence $\mathcal{M}_{1},\dots,\mathcal{M}_{5}$ of margin graphs for the proof, an integer linear program was used to find a small profile $\mathbf{P}_{1}$ and numbers of rankings to add/subtract to keep the total number of voters down while still satisfying all the constraints needed for each step of the proof.
+Then let $\mathbf{P}_{2},\dots,\mathbf{P}_{5}$ be obtained from $\mathbf{P}_{1}$ by adding and subtracting the profiles shown in the equations in Figure 2: e.g., $\mathbf{P}_{2}$ is obtained from $\mathbf{P}_{1}$ by adding 26 new voters who rank $a$ above $d$, $d$ above $b$, etc.; $\mathbf{P}_{3}$ is obtained from $\mathbf{P}_{2}$ by subtracting 77 voters who rank $d$ above $a$, $a$ above $c$, etc.
+Note there are indeed such 77 voters in $\mathbf{P}_{1}$, and for the purposes of $\mathbf{P}_{5}$, there are 77 voters in $\mathbf{P}_{1}$ who rank $d$ above $b$, $b$ above $a$, etc.
 Then $\mathcal{M}_{2},\dots,\mathcal{M}_{5}$ are the margin graphs of $\mathbf{P}_{2},\dots,\mathbf{P}_{5}$, respectively.
@@
 Since $\mathbf{P}_{2}+\mathbf{S}_{2}$ is obtained from $\mathbf{P}_{3}+\mathbf{S}_{2}$ by adding voters all of whom rank $d$ uniquely first, $F(\mathbf{P}_{2}+\mathbf{S}_{2})=\{a\}$} and positive involvement together imply $d\not\in F(\mathbf{P}_{3}+\mathbf{S}_{2})$.
-Since any difference between weights of two distinct edges in $\mathcal{M}_{3}$ is of size at least $4$, no matter which single-voter (or empty) profile $\mathbf{S}_{2}$ is, $b$ is the only defensible candidate besides $d$ in $\mathbf{P}_{3}+\mathbf{S}_{2}$, and $\mathbf{P}_{3}+\mathbf{S}_{2}$ satisfies the margin separation assumption of Lemma 3, so $d\not\in F(\mathbf{P}_{3}+\mathbf{S}_{2})$ implies $F(\mathbf{P}_{3}+\mathbf{S}_{2})=\{b\}$ by Lemma 3.
+Since any difference between weights of two distinct edges in $\mathcal{M}_{3}$ is of size at least 44, no matter which single-voter (or empty) profile $\mathbf{S}_{2}$ is, $b$ is the only defensible candidate besides $d$ in $\mathbf{P}_{3}+\mathbf{S}_{2}$, and $\mathbf{P}_{3}+\mathbf{S}_{2}$ satisfies the margin separation assumption of Lemma 3, so $d\not\in F(\mathbf{P}_{3}+\mathbf{S}_{2})$ implies $F(\mathbf{P}_{3}+\mathbf{S}_{2})=\{b\}$ by Lemma 3.
 Then since $\mathbf{P}_{4}+\mathbf{S}_{2}$ is obtained from $\mathbf{P}_{3}+\mathbf{S}_{2}$ by adding voters all of whom rank $b$ uniquely first, $F(\mathbf{P}_{3}+\mathbf{S}_{2})=\{b\}$ and positive involvement together imply $b\in F(\mathbf{P}_{4}+\mathbf{S}_{2})$.
@@
 Since $\mathbf{P}_{4}+\mathbf{S}_{2}+\mathbf{S}_{4}$ is obtained from $\mathbf{P}_{5}+\mathbf{S}_{2}+\mathbf{S}_{4}$ by adding voters all of whom rank $d$ uniquely first, $F(\mathbf{P}_{4}+\mathbf{S}_{2}+\mathbf{S}_{4})=\{b\}$ and positive involvement together imply $d\not\in F(\mathbf{P}_{5}+\mathbf{S}_{2}+\mathbf{S}_{4})$.
-Finally, since any difference between weights of two distinct edges in $\mathcal{M}_{5}$ is of size at least $6$, no matter which single-voter (or empty) profiles $\mathbf{S}_{2}$ and $\mathbf{S}_{4}$ are, $d$ is the only defensible candidate in $\mathbf{P}_{5}+\mathbf{S}_{2}+\mathbf{S}_{4}$, and $\mathbf{P}_{5}+\mathbf{S}_{2}+\mathbf{S}_{4}$ satisfies the margin separation assumption of Lemma 3, so $F(\mathbf{P}_{5}+\mathbf{S}_{2}+\mathbf{S}_{4})=\{d\}$ by Lemma 3.
+Finally, since any difference between weights of two distinct edges in $\mathcal{M}_{5}$ is of size at least 66, no matter which single-voter (or empty) profiles $\mathbf{S}_{2}$ and $\mathbf{S}_{4}$ are, $d$ is the only defensible candidate in $\mathbf{P}_{5}+\mathbf{S}_{2}+\mathbf{S}_{4}$, and $\mathbf{P}_{5}+\mathbf{S}_{2}+\mathbf{S}_{4}$ satisfies the margin separation assumption of Lemma 3, so $F(\mathbf{P}_{5}+\mathbf{S}_{2}+\mathbf{S}_{4})=\{d\}$ by Lemma 3.
 Therefore, we have a contradiction.∎
@@
 Replace each voter in the profiles $\mathbf{P}_{1},\dots,\mathbf{P}_{5}$ by $n$ copies of that voter, which results in a multiplication of all margins in $\mathcal{M}_{1},\dots,\mathcal{M}_{5}$ by $n$, and then the proof from Theorem 4 goes through as before but now with $|V(\mathbf{S}_{2})|\leq n$ and $|V(\mathbf{S}_{4})|\leq n$.∎
-Thus, for any method satisfying the Condorcet winner and loser criteria and positive involvement, no matter the size of $n$, sometimes a tied winner cannot be made a unique winner even by the addition of up to $n$ voters.
+The foregoing proofs can also be adapted to use negative involvement (cf.
+Fishburn and Brams 1983) in place of positive involvement: for every profile $\mathbf{P}$, $x\in X(\mathbf{P})$, and profile $\mathbf{P}^{\mathord{\downarrow}x}$ with $V(\mathbf{P})\cap V(\mathbf{P}^{\mathord{\downarrow}x})=\varnothing$ and $X(\mathbf{P})=X(\mathbf{P}^{\mathord{\downarrow}x})$, if $x\not\in F(\mathbf{P})$ and $\mathbf{P}^{\mathord{\downarrow}x}$ consists of a single voter who ranks $x$ uniquely last, then $x\not\in F(\mathbf{P}+\mathbf{P}^{\mathord{\downarrow}x})$.
+Contrapositively, if $x\in F(\mathbf{P}+\mathbf{P}^{\mathord{\downarrow}x})$, then $x\in F(\mathbf{P})$.
+Theorem 7.
+There is no voting method satisfying the Condorcet winner and loser criteria, positive or negative involvement, and resolvability.
+Proof.
+The case of positive involvement is given by Theorem 4.
+For negative involvement, first observe that a variant of Lemma 3 holds for negative involvement (cf.
+Pérez 2001, Lemma 1): Let $F$ be a voting method satisfying negative involvement and the Condorcet winner criterion.
+Let $\mathbf{P}$ be a profile in which for all $x,x^{\prime},y,y^{\prime}\in X(\mathbf{P})$ with $x\neq x^{\prime}$ and $y\neq y^{\prime}$, if $\text{Margin}_{\mathbf{P}}(x,x^{\prime})>\text{Margin}_{\mathbf{P}}(y,y^{\prime})$, then $\text{Margin}_{\mathbf{P}}(x,x^{\prime})>\text{Margin}_{\mathbf{P}}(y,y^{\prime})+1$.
+Where $j$ is the largest margin in $\mathcal{M}(\mathbf{P})$, additionally assume that $\mathbf{P}$ contains for each linear order of $X(\mathbf{P})$, $j+1$ voters submitting that linear order.
+[^7] This is a gratuitous assumption that can be weakened in an obvious way, but optimization is not necessary for the proof.
+Then $F(\mathbf{P})\subseteq D(\mathbf{P})$.
+The proof is the same as that of Lemma 3 except that instead of obtaining $\mathbf{P}^{\prime}$ from $\mathbf{P}$ by adding $k+1$ voters who rank $x$ uniquely first and $y$ uniquely second, followed by any linear order of $X(\mathbf{P})\setminus\{x,y\}$, we obtain $\mathbf{P}^{\prime}$ from $\mathbf{P}$ by removing $k+1$ voters who submit any linear order of $X(\mathbf{P})\setminus\{x,y\}$, followed by $y$ in the penultimate position, followed by $x$ uniquely last; these voters exist in $\mathbf{P}$ by our additional assumption on $\mathbf{P}$.
+Then since the proof of Lemma 3 begins with $x\in F(\mathbf{P})$, repeated application of negative involvement in its contrapositive form implies $x\in F(\mathbf{P}^{\prime})$.
+Removing a ranking $L$ has the same effect on margins as adding the reversed ranking $L^{-1}$, so as in the proof of Lemma 3, $y$ is the Condorcet winner in $\mathbf{P}^{\prime}$, yielding a contradiction.
+Now we adapt the proof of Theorem 4 by adding 150 copies of each linear order of $X(\mathbf{P}_{1})$ to $\mathbf{P}_{1}$, which will ensure that the variant of Lemma 3 is applicable when needed in the proof.
+[^8] One can use an integer linear program to synthesize a new starting profile with far fewer voters such that all constraints needed for the proof (and for an optimized version of the variant of Lemma 3) still hold.
+Then instead of adding 26 voters with the ranking adbecadbec to obtain $\mathbf{P}_{2}$, we remove 26 voters with the ranking cebdacebda; instead of removing 7 voters with the ranking dacebdaceb from $\mathbf{P}_{2}$ to obtain $\mathbf{P}_{3}$, we add 7 voters with the ranking becadbecad; and so on for $\mathbf{P}_{4}$ (removing 23 voters with caedbcaedb) and $\mathbf{P}_{5}$ (adding 7 voters with ecabdecabd).
+The margins of $\mathcal{M}_{1},\dots,\mathcal{M}_{5}$ remain as in the proof of Theorem 4.
+Then the rest of the proof is the same as that of Theorem 4 except that we (i) invoke the variant of Lemma 3 for steps involving defensibility and (ii) use the contrapositive form of negative involvement in place of each application of positive involvement.∎
+By a scaling argument exactly analogous to that in the proof of Theorem 6, we can extend the proof of Theorem 7 to establish the following.
+Theorem 8.
+For each positive integer $n$, there is no voting method satisfying the Condorcet winner and loser criteria, positive or negative involvement, and $n$-voter resolvability.
+Thus, for any voting method satisfying the Condorcet winner and loser criteria and positive or negative involvement, no matter the size of $n$, sometimes a tied winner cannot be made a unique winner even by the addition of up to $n$ voters.
 Acknowledgements

### v5

@@
-The incompatibility of the Condorcet winner and loser criteria with positive or negative involvement and resolvability
+The incompatibility of the Condorcet winner and loser criteria with positive involvement and resolvability
 Abstract
@@
 This impossibility theorem holds for any positive integer $n$.
-In addition, positive involvement can be replaced by negative involvement (if a candidate $x$ loses in an initial preference profile, then adding a voter who ranks $x$ uniquely last cannot cause $x$ to win).
-In a previous note, we proved an analogous result assuming an additional axiom of ordinal margin invariance, which we now show is unnecessary for an impossibility theorem, at least if the desired voting method is defined for five-candidate elections.
+It also holds if either the Condorcet loser criterion is replaced by independence of clones or positive involvement is replaced by negative involvement.
 1 Introduction
@@
 In this note, we prove an impossibility theorem illustrating this tradeoff.
-Here the relevant variable-voter axiom is that of positive involvement (see Saari 1995, Pérez 2001, Holliday and Pacuit 2021): if a candidate $x$ wins in an initial preference profile, then adding a voter who ranks $x$ uniquely first cannot cause $x$ to lose (see Definition 1.3).
-Alternatively, we can use the axiom of negative involvement (see Fishburn and Brams 1983, Pérez 2001): if a candidate $x$ loses in an initial preference profile, then adding a voter who ranks $x$ uniquely last cannot cause $x$ to win.
+The relevant variable-voter axiom is that of positive involvement (Saari 1995, Pérez 2001, Holliday and Pacuit 2021): if a candidate $x$ wins in an initial preference profile, then adding a voter who ranks $x$ uniquely first cannot cause $x$ to lose.
+Alternatively, we can use negative involvement (Fishburn and Brams 1983, Pérez 2001): if a candidate $x$ loses in an initial preference profile, then adding a voter who ranks $x$ uniquely last cannot cause $x$ to win.
 According to Pérez (2001, p.
-605), these “may be seen as the minimum to require concerning the coherence in the winning set when new voters are added.” [^1] This seems too strong, however, as a more plausible minimum would be positive-negative involvement: adding a voter who ranks an initial winner $x$ uniquely first and an initial loser $y$ uniquely last cannot cause $x$ to lose and $y$ to win.
-Another weaker condition is bullet-vote positive involvement: adding a voter who ranks an initial winner $x$ uniquely first—and leaves all other candidates unranked or ranks them as tied after xx—cannot cause $x$ to lose.
-On the candidate side, although the famous Condorcet winner and loser criteria (Definition 1.1-2) are not explicitly variable-candidate axioms, they can be motivated by considering candidates dropping out or joining the race: one might think that if $x$ would win (resp.
+605), these “may be seen as the minimum to require concerning the coherence in the winning set when new voters are added.” [^1] However, a weaker requirement is positive-negative involvement: adding a voter who ranks an initial winner $x$ uniquely first and an initial loser $y$ uniquely last cannot cause $x$ to lose and $y$ to win.
+Another weaker requirement is bullet-vote positive involvement: adding a voter who ranks an initial winner $x$ uniquely first—and leaves all other candidates unranked or ranks them as tied after xx—cannot cause $x$ to lose.
+On the candidate side, although the famous Condorcet winner and loser criteria are not explicitly variable-candidate axioms, they can be motivated by considering candidates dropping out or joining the race: one might think that if $x$ would win (resp.
 lose) in every two-candidate sub-election, then $x$ should still win (resp.
 lose) in the full election—or else the election outcome might be seen as too contingent on which other losing candidates happened to stay in the race.
-[^2] As in the case of positive involvement, so too in the case of the Condorcet criteria—especially the Condorcet loser criterion—the motivation mentioned in the main text is not meant to be decisive.
-Finally, for a formalization of the idea that a voting method should not make ties too common, here we use the resolvability axiom that if $x$ is initially tied for winning, then $x$ can be made the unique winner by adding a single new voter (Definition 1.4).
+Instead of the Condorcet loser criterion, we can use the explicitly variable-candidate axiom of independence of clones (Tideman 1987).
+Finally, for a formalization of the idea that a voting method should not make ties too common, we use the resolvability axiom that if $x$ is initially tied for winning, then $x$ can be made the unique winner by adding a single new voter.
 As Tideman (1986, p.
@@
 Given infinite sets $\mathcal{V}$ and $\mathcal{X}$ of possible voters and possible candidates, respectively, a profile $\mathbf{P}$ is a function assigning to each voter $i$ in some nonempty finite set $V(\mathbf{P})\subseteq\mathcal{V}$ a strict weak order $\mathbf{P}(i)$ on some nonempty finite set $X(\mathbf{P})\subseteq\mathcal{X}$ of candidates.
-We call $\mathbf{P}$ linear if for each $i\in V(\mathbf{P})$, $\mathbf{P}(i)$ is a linear order, so there are no ties in $i$’s ranking.
+We call $\mathbf{P}$ linear if for each $i\in V(\mathbf{P})$, $\mathbf{P}(i)$ is a linear order.
 All results to follow continue to hold if all profiles are linear.
-If $\mathbf{P}$ and $\mathbf{Q}$ are profiles with $V(\mathbf{P})\cap V(\mathbf{Q})=\varnothing$ and $X(\mathbf{P})=X(\mathbf{Q})$, then $\mathbf{P}+\mathbf{Q}$ is the profile with $V(\mathbf{P}+\mathbf{Q})=V(\mathbf{P})\cup V(\mathbf{Q})$ defined in the obvious way: if $i\in V(\mathbf{P})$, then $(\mathbf{P}+\mathbf{Q})(i)=\mathbf{P}(i)$, and if $i\in V(\mathbf{Q})$, then $(\mathbf{P}+\mathbf{Q})(i)=\mathbf{Q}(i)$.
+If $\mathbf{P}$ and $\mathbf{Q}$ are profiles with $V(\mathbf{P})\cap V(\mathbf{Q})=\varnothing$ and $X(\mathbf{P})=X(\mathbf{Q})$, then $\mathbf{P}+\mathbf{Q}$ is the profile with $V(\mathbf{P}+\mathbf{Q})=V(\mathbf{P})\cup V(\mathbf{Q})$ such that for $i\in V(\mathbf{P})$, $(\mathbf{P}+\mathbf{Q})(i)=\mathbf{P}(i)$, and for $i\in V(\mathbf{Q})$, $(\mathbf{P}+\mathbf{Q})(i)=\mathbf{Q}(i)$.
 Given $x,y\in X(\mathbf{P})$, the margin of $x$ vs.
@@
 A voting method is a function $F$ assigning to each profile $\mathbf{P}$ a nonempty subset $F(\mathbf{P})\subseteq X(\mathbf{P})$.
-Note that this builds an assumption of “universal domain” into the definition of a voting method.
+This builds an assumption of “universal domain” into the definition of a voting method.
 Definition 1.
@@
 Each of the axioms in Definition 1 is independent of the others.
-The Borda count (Borda 1781) satisfies all of the axioms except the Condorcet winner criterion.
+The Borda count (Borda 1781) satisfies all the axioms except the Condorcet winner criterion.
 Minimax (Simpson 1969, Kramer 1977) satisfies all of them except the Condorcet loser criterion.
 [^3] Minimax does not satisfy resolvability if we restrict its domain to linear profiles (so the added voter must submit a linear order), but the related leximax method (see Brandt et al.
-2025), wherein the winning candidates are those whose vectors of margins, sorted in increasing order, are lexicographically largest, does satisfy resolvability when its domain is restricted (or not restricted) to linear profiles, as well as the Condorcet winner criterion and positive involvement (for all profiles).
+2025)—wherein the winning candidates are those whose vectors of margins, sorted in increasing order, are lexicographically largest—does satisfy resolvability when its domain is restricted (or not restricted) to linear profiles, as well as the Condorcet winner criterion and positive involvement (for all profiles).
 Ranked Pairs (Tideman 1987) satisfies all of them except positive involvement.
 Split Cycle (Holliday and Pacuit 2023) satisfies all of them except resolvability.
-Below we will prove that there is no voting method satisfying all of the axioms in Definition 1.
-The proof involves reasoning about five-candidate profiles.
-If we restrict the domain of voting methods to two or three-candidate profiles, then the axioms are jointly satisfied by the Minimax (Simpson 1969, Kramer 1977) voting method.
-For four candidates, the axioms are jointly inconsistent with an additional axiom of ordinal margin invariance (Holliday 2024), but we leave open whether they are jointly satisfiable without ordinal margin invariance if we only ask for a voting method defined for four-candidate profiles.
-The impossibility proof in the next section makes use of the following concept.
+Below we prove that there is no voting method satisfying all the axioms in Definition 1.
+The proof involves reasoning about profiles with five candidates.
+It extends the strategy of Holliday 2024, which showed that for four candidates, the axioms are jointly inconsistent with an additional axiom of ordinal margin invariance.
+We leave open whether the axioms are jointly satisfiable without ordinal margin invariance if we only ask for a voting method defined on four-candidate profiles.
+If we restrict the domain to two or three-candidate profiles, then the axioms are jointly satisfied by the Minimax voting method.
+The impossibility proof in the next section makes use of the following concept (cf. Kasper et al. 2019).
 Definition 2.
 Given a profile $\mathbf{P}$, the defensible set $D(\mathbf{P})$ of $\mathbf{P}$ is the set of candidates $x\in X(\mathbf{P})$ such that for all $y\in X(\mathbf{P})$, there is a $z\in X(\mathbf{P})$ such that $\text{Margin}_{\mathbf{P}}(z,y)\geq\text{Margin}_{\mathbf{P}}(y,x)$.
-The following lemma is essentially Lemma 3 of Pérez 1995 (cf. Lemma 1 of Pérez 2001), which is in turn an adaptation of a similar result of Moulin 1988, but we include a proof here for convenience.
+The following lemma is essentially Lemma 3 of Pérez 1995 (cf. Lemma 1 of Pérez 2001), which is in turn an adaptation of a similar result of Moulin 1988, but we include a proof for convenience.
 Lemma 3 (Moulin 1988, Pérez 1995).
@@
 Proof.
-Let $F$ be such a method, and suppose $x\in F(\mathbf{P})$.
+Let $F$ be such a method and $x\in F(\mathbf{P})$.
 For contradiction, suppose $x\not\in D(\mathbf{P})$.
@@
 By Debord’s Theorem (Debord 1987), there is a profile $\mathbf{Q}$ whose margin graph is $\mathcal{M}_{1}$.
-Now let $\mathbf{Q}^{+}$ be obtained from $\mathbf{Q}$ by adding, for each ranking of the five candidates, 77 new voters with that ranking; then the margin graph of $\mathbf{Q}^{+}$ is still $\mathcal{M}_{1}$, and $\mathbf{Q}^{+}$ contains all the rankings needed for the rest of the argument, so one may use $\mathbf{Q}^{+}$ in place of $\mathbf{P}_{1}$.
+Now let $\mathbf{Q}^{+}$ be obtained from $\mathbf{Q}$ by adding, for each ranking of the five candidates, $7$ new voters with that ranking; then the margin graph of $\mathbf{Q}^{+}$ is still $\mathcal{M}_{1}$, and $\mathbf{Q}^{+}$ contains all the rankings needed for the rest of the argument, so one may use $\mathbf{Q}^{+}$ in place of $\mathbf{P}_{1}$.
 After the initial construction of this proof using another version of $\mathcal{M}_{1}$ and different numbers of rankings to add/subtract to obtain an appropriate sequence $\mathcal{M}_{1},\dots,\mathcal{M}_{5}$ of margin graphs for the proof, an integer linear program was used to find a small profile $\mathbf{P}_{1}$ and numbers of rankings to add/subtract to keep the total number of voters down while still satisfying all the constraints needed for each step of the proof.
-Then let $\mathbf{P}_{2},\dots,\mathbf{P}_{5}$ be obtained from $\mathbf{P}_{1}$ by adding and subtracting the profiles shown in the equations in Figure 2: e.g., $\mathbf{P}_{2}$ is obtained from $\mathbf{P}_{1}$ by adding 26 new voters who rank $a$ above $d$, $d$ above $b$, etc.; $\mathbf{P}_{3}$ is obtained from $\mathbf{P}_{2}$ by subtracting 77 voters who rank $d$ above $a$, $a$ above $c$, etc.
-Note there are indeed such 77 voters in $\mathbf{P}_{1}$, and for the purposes of $\mathbf{P}_{5}$, there are 77 voters in $\mathbf{P}_{1}$ who rank $d$ above $b$, $b$ above $a$, etc.
+Even fewer voters suffice if, at the expense of elegance, we allow more distinct ranking types in the profiles.
+Let $\mathbf{P}_{2},\dots,\mathbf{P}_{5}$ be obtained from $\mathbf{P}_{1}$ by adding and subtracting the profiles shown in the equations in Figure 2: $\mathbf{P}_{2}$ is obtained from $\mathbf{P}_{1}$ by adding 26 new voters with the ranking adbecadbec; $\mathbf{P}_{3}$ is obtained from $\mathbf{P}_{2}$ by subtracting $7$ voters with the ranking dacebdaceb, etc.
+Note there are indeed such $7$ voters in $\mathbf{P}_{1}$, and for the purposes of $\mathbf{P}_{5}$, there are $7$ voters in $\mathbf{P}_{1}$ with the ranking dbacedbace.
 Then $\mathcal{M}_{2},\dots,\mathcal{M}_{5}$ are the margin graphs of $\mathbf{P}_{2},\dots,\mathbf{P}_{5}$, respectively.
@@
 Since $\mathbf{P}_{2}+\mathbf{S}_{2}$ is obtained from $\mathbf{P}_{3}+\mathbf{S}_{2}$ by adding voters all of whom rank $d$ uniquely first, $F(\mathbf{P}_{2}+\mathbf{S}_{2})=\{a\}$} and positive involvement together imply $d\not\in F(\mathbf{P}_{3}+\mathbf{S}_{2})$.
-Since any difference between weights of two distinct edges in $\mathcal{M}_{3}$ is of size at least 44, no matter which single-voter (or empty) profile $\mathbf{S}_{2}$ is, $b$ is the only defensible candidate besides $d$ in $\mathbf{P}_{3}+\mathbf{S}_{2}$, and $\mathbf{P}_{3}+\mathbf{S}_{2}$ satisfies the margin separation assumption of Lemma 3, so $d\not\in F(\mathbf{P}_{3}+\mathbf{S}_{2})$ implies $F(\mathbf{P}_{3}+\mathbf{S}_{2})=\{b\}$ by Lemma 3.
+Since any difference between weights of two distinct edges in $\mathcal{M}_{3}$ is of size at least $4$, no matter which single-voter (or empty) profile $\mathbf{S}_{2}$ is, $b$ is the only defensible candidate besides $d$ in $\mathbf{P}_{3}+\mathbf{S}_{2}$, and $\mathbf{P}_{3}+\mathbf{S}_{2}$ satisfies the margin separation assumption of Lemma 3, so $d\not\in F(\mathbf{P}_{3}+\mathbf{S}_{2})$ implies $F(\mathbf{P}_{3}+\mathbf{S}_{2})=\{b\}$ by Lemma 3.
 Then since $\mathbf{P}_{4}+\mathbf{S}_{2}$ is obtained from $\mathbf{P}_{3}+\mathbf{S}_{2}$ by adding voters all of whom rank $b$ uniquely first, $F(\mathbf{P}_{3}+\mathbf{S}_{2})=\{b\}$ and positive involvement together imply $b\in F(\mathbf{P}_{4}+\mathbf{S}_{2})$.
@@
 Since $\mathbf{P}_{4}+\mathbf{S}_{2}+\mathbf{S}_{4}$ is obtained from $\mathbf{P}_{5}+\mathbf{S}_{2}+\mathbf{S}_{4}$ by adding voters all of whom rank $d$ uniquely first, $F(\mathbf{P}_{4}+\mathbf{S}_{2}+\mathbf{S}_{4})=\{b\}$ and positive involvement together imply $d\not\in F(\mathbf{P}_{5}+\mathbf{S}_{2}+\mathbf{S}_{4})$.
-Finally, since any difference between weights of two distinct edges in $\mathcal{M}_{5}$ is of size at least 66, no matter which single-voter (or empty) profiles $\mathbf{S}_{2}$ and $\mathbf{S}_{4}$ are, $d$ is the only defensible candidate in $\mathbf{P}_{5}+\mathbf{S}_{2}+\mathbf{S}_{4}$, and $\mathbf{P}_{5}+\mathbf{S}_{2}+\mathbf{S}_{4}$ satisfies the margin separation assumption of Lemma 3, so $F(\mathbf{P}_{5}+\mathbf{S}_{2}+\mathbf{S}_{4})=\{d\}$ by Lemma 3.
+Finally, since any difference between weights of two distinct edges in $\mathcal{M}_{5}$ is of size at least $6$, no matter which single-voter (or empty) profiles $\mathbf{S}_{2}$ and $\mathbf{S}_{4}$ are, $d$ is the only defensible candidate in $\mathbf{P}_{5}+\mathbf{S}_{2}+\mathbf{S}_{4}$, and $\mathbf{P}_{5}+\mathbf{S}_{2}+\mathbf{S}_{4}$ satisfies the margin separation assumption of Lemma 3, so $F(\mathbf{P}_{5}+\mathbf{S}_{2}+\mathbf{S}_{4})=\{d\}$ by Lemma 3.
 Therefore, we have a contradiction.∎
@@
 Inspection of the proof of Theorem 4 shows that resolvability can be replaced by quasi-resoluteness (in the terminology of Holliday and Pacuit 2023): $|F(\mathbf{P})|=1$ for any profile $\mathbf{P}$ that is uniquely weighted in the sense that for all $x,y,x^{\prime},y^{\prime}\in X(\mathbf{P})$ with $x\neq y$ and $x^{\prime}\neq y^{\prime}$, if $(x,y)\neq(x^{\prime},y^{\prime})$, then $\text{Margin}_{\mathbf{P}}(x,y)\neq\text{Margin}_{\mathbf{P}}(x^{\prime},y^{\prime})$.
-It also shows that the package of positive involvement and resolvability can be replaced by strict positive involvement (in the terminology of Dominik Peters, personal communication), which is defined like positive involvement but with $F(\mathbf{P}+\mathbf{P}^{x})=\{x\}$ in place of $x\in F(\mathbf{P}+\mathbf{P}^{x})$.
+It also shows that the package of positive involvement and resolvability can be replaced by strict positive involvement (in the terminology of Dominik Peters, personal communication), which is defined like positive involvement but with $F(\mathbf{P}+\mathbf{P}^{x})=\{x\}$ instead of $x\in F(\mathbf{P}+\mathbf{P}^{x})$.
 The same proof strategy can be used to show that for each positive integer $n$, resolvability can be replaced by $n$-voter resolvability: for every profile $\mathbf{P}$, if $|F(\mathbf{P})|>1$, then for each $x\in F(\mathbf{P})$, there is a profile $\mathbf{Q}$ with $|V(\mathbf{Q})|\leq n$, $V(\mathbf{P})\cap V(\mathbf{Q})=\varnothing$, and $X(\mathbf{P})=X(\mathbf{Q})$ such that $F(\mathbf{P}+\mathbf{Q})=\{x\}$.
@@
 Proof.
-Replace each voter in the profiles $\mathbf{P}_{1},\dots,\mathbf{P}_{5}$ by $n$ copies of that voter, which results in a multiplication of all margins in $\mathcal{M}_{1},\dots,\mathcal{M}_{5}$ by $n$, and then the proof from Theorem 4 goes through as before but now with $|V(\mathbf{S}_{2})|\leq n$ and $|V(\mathbf{S}_{4})|\leq n$.∎
-The foregoing proofs can also be adapted to use negative involvement (cf.
-Fishburn and Brams 1983) in place of positive involvement: for every profile $\mathbf{P}$, $x\in X(\mathbf{P})$, and profile $\mathbf{P}^{\mathord{\downarrow}x}$ with $V(\mathbf{P})\cap V(\mathbf{P}^{\mathord{\downarrow}x})=\varnothing$ and $X(\mathbf{P})=X(\mathbf{P}^{\mathord{\downarrow}x})$, if $x\not\in F(\mathbf{P})$ and $\mathbf{P}^{\mathord{\downarrow}x}$ consists of a single voter who ranks $x$ uniquely last, then $x\not\in F(\mathbf{P}+\mathbf{P}^{\mathord{\downarrow}x})$.
+Replace each voter in $\mathbf{P}_{1},\dots,\mathbf{P}_{5}$ by $n$ copies of that voter, which results in a multiplication of each margin in $\mathcal{M}_{1},\dots,\mathcal{M}_{5}$ by $n$, and then the proof from Theorem 4 goes through as before but now with $|V(\mathbf{S}_{2})|\leq n$ and $|V(\mathbf{S}_{4})|\leq n$.∎
+These proofs can also be adapted to use negative involvement (Fishburn and Brams 1983) instead of positive involvement: for every profile $\mathbf{P}$, $x\in X(\mathbf{P})$, and profile $\mathbf{P}^{\mathord{\downarrow}x}$ with $V(\mathbf{P})\cap V(\mathbf{P}^{\mathord{\downarrow}x})=\varnothing$ and $X(\mathbf{P})=X(\mathbf{P}^{\mathord{\downarrow}x})$, if $x\not\in F(\mathbf{P})$ and $\mathbf{P}^{\mathord{\downarrow}x}$ consists of a single voter who ranks $x$ uniquely last, then $x\not\in F(\mathbf{P}+\mathbf{P}^{\mathord{\downarrow}x})$.
 Contrapositively, if $x\in F(\mathbf{P}+\mathbf{P}^{\mathord{\downarrow}x})$, then $x\in F(\mathbf{P})$.
 Theorem 7.
-There is no voting method satisfying the Condorcet winner and loser criteria, positive or negative involvement, and resolvability.
+There is no voting method satisfying the Condorcet winner and loser criteria, negative involvement, and resolvability.
 Proof.
-The case of positive involvement is given by Theorem 4.
-For negative involvement, first observe that a variant of Lemma 3 holds for negative involvement (cf.
+A variant of Lemma 3 holds for negative involvement (cf.
 Pérez 2001, Lemma 1): Let $F$ be a voting method satisfying negative involvement and the Condorcet winner criterion.
 Let $\mathbf{P}$ be a profile in which for all $x,x^{\prime},y,y^{\prime}\in X(\mathbf{P})$ with $x\neq x^{\prime}$ and $y\neq y^{\prime}$, if $\text{Margin}_{\mathbf{P}}(x,x^{\prime})>\text{Margin}_{\mathbf{P}}(y,y^{\prime})$, then $\text{Margin}_{\mathbf{P}}(x,x^{\prime})>\text{Margin}_{\mathbf{P}}(y,y^{\prime})+1$.
-Where $j$ is the largest margin in $\mathcal{M}(\mathbf{P})$, additionally assume that $\mathbf{P}$ contains for each linear order of $X(\mathbf{P})$, $j+1$ voters submitting that linear order.
-[^7] This is a gratuitous assumption that can be weakened in an obvious way, but optimization is not necessary for the proof.
+Additionally assume that for each $x\in X(\mathbf{P})\setminus D(\mathbf{P})$, there is a $y\in X(\mathbf{P})$ such that for every $z\in X(\mathbf{P})$, $\text{Margin}_{\mathbf{P}}(z,y)<\text{Margin}_{\mathbf{P}}(y,x)$, and where $k=\max\{\text{Margin}_{\mathbf{P}}(z,y)\mid z\in X(\mathbf{P}),z\neq y\}$, there are at least $k+1$ voters whose rankings have $y$ uniquely second-to-last, followed by $x$ uniquely last.
 Then $F(\mathbf{P})\subseteq D(\mathbf{P})$.
-The proof is the same as that of Lemma 3 except that instead of obtaining $\mathbf{P}^{\prime}$ from $\mathbf{P}$ by adding $k+1$ voters who rank $x$ uniquely first and $y$ uniquely second, followed by any linear order of $X(\mathbf{P})\setminus\{x,y\}$, we obtain $\mathbf{P}^{\prime}$ from $\mathbf{P}$ by removing $k+1$ voters who submit any linear order of $X(\mathbf{P})\setminus\{x,y\}$, followed by $y$ in the penultimate position, followed by $x$ uniquely last; these voters exist in $\mathbf{P}$ by our additional assumption on $\mathbf{P}$.
+The proof is the same as that of Lemma 3 except we obtain $\mathbf{P}^{\prime}$ from $\mathbf{P}$ by removing $k+1$ voters of the kind described above.
 Then since the proof of Lemma 3 begins with $x\in F(\mathbf{P})$, repeated application of negative involvement in its contrapositive form implies $x\in F(\mathbf{P}^{\prime})$.
 Removing a ranking $L$ has the same effect on margins as adding the reversed ranking $L^{-1}$, so as in the proof of Lemma 3, $y$ is the Condorcet winner in $\mathbf{P}^{\prime}$, yielding a contradiction.
-Now we adapt the proof of Theorem 4 by adding 150 copies of each linear order of $X(\mathbf{P}_{1})$ to $\mathbf{P}_{1}$, which will ensure that the variant of Lemma 3 is applicable when needed in the proof.
-[^8] One can use an integer linear program to synthesize a new starting profile with far fewer voters such that all constraints needed for the proof (and for an optimized version of the variant of Lemma 3) still hold.
-Then instead of adding 26 voters with the ranking adbecadbec to obtain $\mathbf{P}_{2}$, we remove 26 voters with the ranking cebdacebda; instead of removing 7 voters with the ranking dacebdaceb from $\mathbf{P}_{2}$ to obtain $\mathbf{P}_{3}$, we add 7 voters with the ranking becadbecad; and so on for $\mathbf{P}_{4}$ (removing 23 voters with caedbcaedb) and $\mathbf{P}_{5}$ (adding 7 voters with ecabdecabd).
-The margins of $\mathcal{M}_{1},\dots,\mathcal{M}_{5}$ remain as in the proof of Theorem 4.
-Then the rest of the proof is the same as that of Theorem 4 except that we (i) invoke the variant of Lemma 3 for steps involving defensibility and (ii) use the contrapositive form of negative involvement in place of each application of positive involvement.∎
+Now we adapt the proof of Theorem 4 by adding 147 copies of each linear order of $X(\mathbf{P}_{1})$ to $\mathbf{P}_{1}$, which ensures that the variant of Lemma 3 is applicable when needed in the proof.
+[^6] By starting with at least $147$ copies of each linear order in $\mathbf{P}_{1}$, we ensure that there are at least $121$ copies of each linear order in $\mathbf{P}_{3}+\mathbf{S}_{2}$ and in $\mathbf{P}_{5}+\mathbf{S}_{2}+\mathbf{S}_{4}$, which is greater than the maximal margins in those profiles, so the variant of Lemma 3 is applicable to those profiles.
+Of course, one can use an integer linear program to synthesize a new starting profile with far fewer voters such that all constraints needed for the proof still hold.
+Then instead of adding 26 voters with the ranking adbecadbec to obtain $\mathbf{P}_{2}$, we remove 26 voters with cebdacebda; instead of removing 7 voters with dacebdaceb from $\mathbf{P}_{2}$ to obtain $\mathbf{P}_{3}$, we add 7 voters with becadbecad; and so on for $\mathbf{P}_{4}$ (removing 23 voters with caedbcaedb) and $\mathbf{P}_{5}$ (adding 7 voters with ecabdecabd).
+The margins of $\mathcal{M}_{1},\dots,\mathcal{M}_{5}$ remain as in the proof of Theorem 4, and then the rest of the proof is the same except that we (i) invoke the variant of Lemma 3 for steps involving defensibility and (ii) use negative involvement instead of positive involvement.∎
 By a scaling argument exactly analogous to that in the proof of Theorem 6, we can extend the proof of Theorem 7 to establish the following.
 Theorem 8.
-For each positive integer $n$, there is no voting method satisfying the Condorcet winner and loser criteria, positive or negative involvement, and $n$-voter resolvability.
-Thus, for any voting method satisfying the Condorcet winner and loser criteria and positive or negative involvement, no matter the size of $n$, sometimes a tied winner cannot be made a unique winner even by the addition of up to $n$ voters.
+For each positive integer $n$, there is no voting method satisfying the Condorcet winner and loser criteria, negative involvement, and $n$-voter resolvability.
+Finally, in response to a previous version of this note with the results above, Dominik Peters (personal communication) suggested that analogous impossibilities might hold with independence of clones (Tideman 1987) instead of the Condorcet loser criterion.
+Indeed, we can confirm that such a swap in Theorems 4 and 6 is possible, as well as in Theorems 78 if we add the axiom of block preservation (Holliday and Pacuit 2025): if the profile $\mathbf{L}$ consists of exactly one copy of each linear order of $X(\mathbf{P})$, then $F(\mathbf{P})\subseteq F(\mathbf{P}+\mathbf{L})$.
+Recall that a set $C\subsetneq X(\mathbf{P})$ with $|C|\geq 2$ is a set of clones in $\mathbf{P}$ if for each $i\in V(\mathbf{P})$ and $x\in X(\mathbf{P})\setminus C$, either $i$ ranks $x$ above each $c\in C$ or $i$ ranks $x$ below each $c\in C$.
+Given a profile $\mathbf{P}$ and $c\in X(\mathbf{P})$, let $\mathbf{P}_{-c}$ be the profile with $X(\mathbf{P}_{-c})=X(\mathbf{P})\setminus\{c\}$ and $V(\mathbf{P}_{-c})=V(\mathbf{P})$ where for each $i\in V(\mathbf{P})$, $\mathbf{P}_{-c}(i)$ is the restriction of $\mathbf{P}(i)$ to $X(\mathbf{P}_{-c})$.
+Then $F$ satisfies independence of clones if for any profile $\mathbf{P}$, set $C$ of clones, $c\in C$, and $x\in X(\mathbf{P})\setminus C$, (i) $x\in F(\mathbf{P})$ iff $x\in F(\mathbf{P}_{-c})$, and (ii) $C\cap F(\mathbf{P})\neq\varnothing$ iff $C\cap F(\mathbf{P}_{-c})\neq\varnothing$.
+Theorem 9.
+For each positive integer $n$, there is no voting method satisfying the Condorcet winner criterion, independence of clones, positive involvement, and $n$-voter resolvability.
+The same holds with positive involvement replaced by negative involvement plus block preservation.
+Proof.
+We prove the $n=1$ case.
+Reasoning as in the proof of Theorem 6 handles $n>1$.
+For positive involvement, consider a $\mathbf{Q}_{1}$ as in Figure 3, wherein $\{a,b,c,e\}$ is a set of clones.
+By Lemma 3, $F(\mathbf{Q}_{1})\subseteq\{a,d\}$.
+By the Condorcet winner criterion, $F(\mathbf{Q}_{1-b})=\{a\}$, so by independence of clones (part (i) or (ii)) and $F(\mathbf{Q}_{1})\subseteq\{a,d\}$, we have $a\in F(\mathbf{Q}_{1})$.
+Then obtain $\mathbf{Q}_{2},\dots,\mathbf{Q}_{5}$ from $\mathbf{Q}_{1}$ by adding and subtracting the profiles in Figure 4.
+The rest of the proof is just like that of Theorem 4 but with $\mathbf{Q}_{i}$ instead of $\mathbf{P}_{i}$.
+For negative involvement, by block preservation, we can add as many blocks of all linear orders as we wish to $\mathbf{Q}_{1}$ to obtain a $\mathbf{Q}_{1}^{+}$ with $F(\mathbf{Q}_{1})\subseteq F(\mathbf{Q}_{1}^{+})\subseteq\{a,d\}$ by the variant of Lemma 3 in the proof of Theorem 7.
+Hence $F(\mathbf{Q}_{1})\subseteq\{a,d\}$, which implies $a\in F(\mathbf{Q}_{1})$ as above and therefore $a\in F(\mathbf{Q}_{1}^{+})$.
+Then starting from $\mathbf{Q}_{1}^{+}$, we reason with negative involvement instead of positive involvement as in the proof of Theorem 7.∎
+62 60 42 23 19 3
+$d$ $c$ $d$ $a$ $e$ $c$
+$b$ $b$ $e$ $e$ $c$ $e$
+$a$ $a$ $a$ $c$ $b$ $b$
+$e$ $e$ $c$ $b$ $a$ $a$
+$c$ $d$ $b$ $d$ $d$ $d$
+Figure 3: the profile $\mathbf{Q}_{1}$ used in the proof of Theorem 9.
+$\mathcal{M}_{1}$
+aabbccddee45451181817979113535858511838311
++\;+ 16 7 2 $a$ $a$ $a$ $d$ $e$ $e$ $b$ $d$ $b$ $c$ $b$ $d$ $e$ $c$ $c$ ==\;\; $\mathcal{M}_{2}$aabbccddee7070262610610654544242606020202424667676 == 5 $d$ $e$ $a$ $c$ $b$ ++\;\; $\mathcal{M}_{3}$aabbccddee6565313111111159594747555515151919117171
+$\mathcal{M}_{3}$
+aabbccddee6565313111111159594747555515151919117171
++\,\;\;\;\,+\;\;\; 20 $b$ $d$ $c$ $a$ $e$ =\;\;\;=\;\;\;\;\;\,\, $\mathcal{M}_{4}$aabbccddee4545111113113179795567673535393921215151 == 11 $d$ $b$ $a$ $e$ $c$ ++\;\; $\mathcal{M}_{5}$aabbccddee343422221201206868161656564646282810104040
+Figure 4: margin graphs used in the proof of Theorem 9 with defensible candidates shaded gray.
 Acknowledgements
-I thank Emma Boniface and Diego Bejarano for helpful discussions of the problem posed at the end of Holliday 2024 and Dominik Peters for helpful feedback on this paper.
+I thank Diego Bejarano and Emma Boniface for helpful discussions of the problem posed at the end of Holliday 2024 and Dominik Peters for helpful feedback on this paper.

### v6

@@
 Another weaker requirement is bullet-vote positive involvement: adding a voter who ranks an initial winner $x$ uniquely first—and leaves all other candidates unranked or ranks them as tied after xx—cannot cause $x$ to lose.
-On the candidate side, although the famous Condorcet winner and loser criteria are not explicitly variable-candidate axioms, they can be motivated by considering candidates dropping out or joining the race: one might think that if $x$ would win (resp.
+On the candidate side, although the famous Condorcet winner and loser criteria (Condorcet 1785) are not explicitly variable-candidate axioms, they can be motivated by considering candidates dropping out or joining the race: one might think that if $x$ would win (resp.
 lose) in every two-candidate sub-election, then $x$ should still win (resp.
@@
 Instead of the Condorcet loser criterion, we can use the explicitly variable-candidate axiom of independence of clones (Tideman 1987).
-Finally, for a formalization of the idea that a voting method should not make ties too common, we use the resolvability axiom that if $x$ is initially tied for winning, then $x$ can be made the unique winner by adding a single new voter.
+Finally, for a formalization of the idea that a voting method should not make ties too common, we use the resolvability axiom (Tideman 1986) that if $x$ is initially tied for winning, then $x$ can be made the unique winner by adding a single new voter.
 As Tideman (1986, p.
-25) puts it, “Resolvability may be thought of as a condition that ties happen only “by coincidence”, every tie being within one vote of being broken.” In fact, we will ultimately consider, for any positive integer $n$, the axiom of $n$-voter resolvability, whereby any tied winner can be made the unique winner by adding some set of up to $n$ new voters.
+25) puts it, “Resolvability may be thought of as a condition that ties happen only ‘by coincidence’, every tie being within one vote of being broken.” In fact, we will ultimately consider, for any positive integer $n$, the axiom of $n$-voter resolvability, whereby any tied winner can be made the unique winner by adding some set of up to $n$ new voters.
 2 Setup
@@
   1. 1.
-    $F$ satisfies the Condorcet winner criterion (Condorcet 1785) if for every profile $\mathbf{P}$ and $x\in X(\mathbf{P})$, if $\text{Margin}_{\mathbf{P}}(x,y)>0$ for all $y\in X(\mathbf{P})\setminus\{x\}$ (so $x$ is a Condorcet winner), then $F(\mathbf{P})=\{x\}$.
+    $F$ satisfies the Condorcet winner criterion if for every profile $\mathbf{P}$ and $x\in X(\mathbf{P})$, if $\text{Margin}_{\mathbf{P}}(x,y)>0$ for all $y\in X(\mathbf{P})\setminus\{x\}$ (so $x$ is a Condorcet winner), then $F(\mathbf{P})=\{x\}$.
   2. 2.
-    $F$ satisfies the Condorcet loser criterion (Condorcet 1785) if for every profile $\mathbf{P}$ and $x\in X(\mathbf{P})$, if $\text{Margin}_{\mathbf{P}}(y,x)>0$ for all $y\in X(\mathbf{P})\setminus\{x\}$ (so $x$ is a Condorcet loser), then $x\not\in F(\mathbf{P})$.
+    $F$ satisfies the Condorcet loser criterion if for every profile $\mathbf{P}$ and $x\in X(\mathbf{P})$, if $\text{Margin}_{\mathbf{P}}(y,x)>0$ for all $y\in X(\mathbf{P})\setminus\{x\}$ (so $x$ is a Condorcet loser), then $x\not\in F(\mathbf{P})$.
   3. 3.
-    $F$ satisfies positive involvement (Saari 1995) if for every profile $\mathbf{P}$, $x\in X(\mathbf{P})$, and profile $\mathbf{P}^{x}$ with $V(\mathbf{P})\cap V(\mathbf{P}^{x})=\varnothing$ and $X(\mathbf{P})=X(\mathbf{P}^{x})$, if $x\in F(\mathbf{P})$ and $\mathbf{P}^{x}$ consists of a single voter who ranks $x$ uniquely first, then $x\in F(\mathbf{P}+\mathbf{P}^{x})$.
+    $F$ satisfies positive involvement if for every profile $\mathbf{P}$, $x\in X(\mathbf{P})$, and profile $\mathbf{P}^{x}$ with $V(\mathbf{P})\cap V(\mathbf{P}^{x})=\varnothing$ and $X(\mathbf{P})=X(\mathbf{P}^{x})$, if $x\in F(\mathbf{P})$ and $\mathbf{P}^{x}$ consists of a single voter who ranks $x$ uniquely first, then $x\in F(\mathbf{P}+\mathbf{P}^{x})$.
   4. 4.
-    $F$ satisfies resolvability (Tideman 1986) if for every profile $\mathbf{P}$, if $|F(\mathbf{P})|>1$, then for each $x\in F(\mathbf{P})$, there is some single-voter profile $\mathbf{S}$ with $V(\mathbf{P})\cap V(\mathbf{S})=\varnothing$ and $X(\mathbf{P})=X(\mathbf{S})$ such that $F(\mathbf{P}+\mathbf{S})=\{x\}$.
-[^2] In Holliday 2024, we used a weaker version of this axiom in the impossibility theorem, since there the weaker version sufficed: if $|F(\mathbf{P})|>1$, then there is some single-voter profile $\mathbf{S}$ such $|F(\mathbf{P}+\mathbf{S})|=1$.
-But Tideman’s original axiom (here stated in a variable-voter setting) requires that if $|F(\mathbf{P})|>1$, then for each $x\in F(\mathbf{P})$, there is some single-voter profile $\mathbf{S}$ such that $F(\mathbf{P}+\mathbf{S})=\{x\}$.
-This version facilitates the proof in this paper.
+    $F$ satisfies resolvability if for every profile $\mathbf{P}$, if $|F(\mathbf{P})|>1$, then for each $x\in F(\mathbf{P})$, there is some single-voter profile $\mathbf{S}$ with $V(\mathbf{P})\cap V(\mathbf{S})=\varnothing$ and $X(\mathbf{P})=X(\mathbf{S})$ such that $F(\mathbf{P}+\mathbf{S})=\{x\}$.
+[^2] In Holliday 2024, we used a weaker version of this axiom in the impossibility theorem, since there the weaker version sufficed: if $|F(\mathbf{P})|>1$, then there is some single-voter profile $\mathbf{S}$ such that $|F(\mathbf{P}+\mathbf{S})|=1$.
+But Tideman’s (1986) original axiom requires that for each $x\in F(\mathbf{P})$, there is some single-voter profile $\mathbf{S}$ such that $F(\mathbf{P}+\mathbf{S})=\{x\}$.
+This version facilitates the proof in this paper (Tideman 1986 does not include the assumption that $|F(\mathbf{P})|>1$, but the language of Tideman 1987 suggests it).
 In Holliday 2024, we also discussed asymptotic resolvability, but without the assumption of ordinal margin invariance made in that paper, asymptotic resolvability is more difficult to work with, so we do not discuss it here.
 Each of the axioms in Definition 1 is independent of the others.
-The Borda count (Borda 1781) satisfies all the axioms except the Condorcet winner criterion.
+The Borda count (Borda 1784) satisfies all the axioms except the Condorcet winner criterion.
 Minimax (Simpson 1969, Kramer 1977) satisfies all of them except the Condorcet loser criterion.
@@
 We leave open whether the axioms are jointly satisfiable without ordinal margin invariance if we only ask for a voting method defined on four-candidate profiles.
-If we restrict the domain to two or three-candidate profiles, then the axioms are jointly satisfied by the Minimax voting method.
+If we restrict the domain to two- or three-candidate profiles, then the axioms are jointly satisfied by the Minimax voting method.
 The impossibility proof in the next section makes use of the following concept (cf. Kasper et al. 2019).
@@
 After the initial construction of this proof using another version of $\mathcal{M}_{1}$ and different numbers of rankings to add/subtract to obtain an appropriate sequence $\mathcal{M}_{1},\dots,\mathcal{M}_{5}$ of margin graphs for the proof, an integer linear program was used to find a small profile $\mathbf{P}_{1}$ and numbers of rankings to add/subtract to keep the total number of voters down while still satisfying all the constraints needed for each step of the proof.
-Even fewer voters suffice if, at the expense of elegance, we allow more distinct ranking types in the profiles.
+Even fewer voters suffice (e.g., $247$ instead of $261$ in $\mathbf{P}_{4}$) if, at the expense of elegance, we allow more distinct ranking types in the profiles.
 Let $\mathbf{P}_{2},\dots,\mathbf{P}_{5}$ be obtained from $\mathbf{P}_{1}$ by adding and subtracting the profiles shown in the equations in Figure 2: $\mathbf{P}_{2}$ is obtained from $\mathbf{P}_{1}$ by adding 26 new voters with the ranking adbecadbec; $\mathbf{P}_{3}$ is obtained from $\mathbf{P}_{2}$ by subtracting $7$ voters with the ranking dacebdaceb, etc.
@@
 Inspection of the proof of Theorem 4 shows that resolvability can be replaced by quasi-resoluteness (in the terminology of Holliday and Pacuit 2023): $|F(\mathbf{P})|=1$ for any profile $\mathbf{P}$ that is uniquely weighted in the sense that for all $x,y,x^{\prime},y^{\prime}\in X(\mathbf{P})$ with $x\neq y$ and $x^{\prime}\neq y^{\prime}$, if $(x,y)\neq(x^{\prime},y^{\prime})$, then $\text{Margin}_{\mathbf{P}}(x,y)\neq\text{Margin}_{\mathbf{P}}(x^{\prime},y^{\prime})$.
-It also shows that the package of positive involvement and resolvability can be replaced by strict positive involvement (in the terminology of Dominik Peters, personal communication), which is defined like positive involvement but with $F(\mathbf{P}+\mathbf{P}^{x})=\{x\}$ instead of $x\in F(\mathbf{P}+\mathbf{P}^{x})$.
+In addition, the package of positive involvement and resolvability can be replaced by strict positive involvement (in the terminology of Dominik Peters, personal communication), which is defined like positive involvement but with $F(\mathbf{P}+\mathbf{P}^{x})=\{x\}$ instead of $x\in F(\mathbf{P}+\mathbf{P}^{x})$.
+[^6] A slight modification of the proof of Theorem 4 shows that positive involvement can be replaced by the weaker axiom that if $F(\mathbf{P})=\{x\}$ and $\mathbf{P}^{x}$ is a profile of new voters all of whom rank $x$ uniquely first, then $x\in F(\mathbf{P}+\mathbf{P}^{x})$ (cf.
+Richelson’s (1978) voter adaptability, which requires that $F(\mathbf{P})=\{x\}$ implies $F(\mathbf{P}+\mathbf{P}^{x})=\{x\}$).
+First, a variant of Lemma 3 holds with the weaker axiom in place of positive involvement, provided we assume $F$ satisfies resolvability and that $\text{Margin}_{\mathbf{P}}(x,x^{\prime})>\text{Margin}_{\mathbf{P}}(y,y^{\prime})$ implies $\text{Margin}_{\mathbf{P}}(x,x^{\prime})>\text{Margin}_{\mathbf{P}}(y,y^{\prime})+3$.
+Now double each voter in each profile in the proof of Theorem 4, so all margins double.
+With the weaker axiom, we derive $F(\mathbf{P}_{3}+\mathbf{S}_{2})\neq\{d\}$ and then $b\in F(\mathbf{P}_{3}+\mathbf{S}_{2})$ by the variant of Lemma 3.
+Then by resolvability, we add an “empty” or single-voter $\mathbf{S}_{3}$ such that $F(\mathbf{P}_{3}+\mathbf{S}_{2}+\mathbf{S}_{3})=\{b\}$.
+From here the proof proceeds as before but to the conclusion that $F(\mathbf{P}_{5}+\mathbf{S}_{2}+\mathbf{S}_{3}+\mathbf{S}_{4})\neq\{d\}$.
+By the variant of Lemma 3, which is still applicable due to the doubling of margins, we have $F(\mathbf{P}_{5}+\mathbf{S}_{2}+\mathbf{S}_{3}+\mathbf{S}_{4})=\{d\}$.
+Similarly, the negative involvement axiom in Theorem 7 can be replaced by the weaker axiom that if $x\not\in F(\mathbf{P})$ and $\mathbf{P}^{\downarrow x}$ is a profile of new voters all of whom rank $x$ uniquely last, then $F(\mathbf{P}+\mathbf{P}^{\downarrow x})\neq\{x\}$ (cf.
+Brandt et al.’s (2025) singleton negative involvement, which requires that $F(\mathbf{P})\neq\{x\}$ implies $F(\mathbf{P}+\mathbf{P}^{\downarrow x})\neq\{x\}$).
 The same proof strategy can be used to show that for each positive integer $n$, resolvability can be replaced by $n$-voter resolvability: for every profile $\mathbf{P}$, if $|F(\mathbf{P})|>1$, then for each $x\in F(\mathbf{P})$, there is a profile $\mathbf{Q}$ with $|V(\mathbf{Q})|\leq n$, $V(\mathbf{P})\cap V(\mathbf{Q})=\varnothing$, and $X(\mathbf{P})=X(\mathbf{Q})$ such that $F(\mathbf{P}+\mathbf{Q})=\{x\}$.
+[^7] We could instead consider $r$-resolvability: for some $N$ and every $\mathbf{P}$ with $|V(\mathbf{P})|\geq N$, if $|F(\mathbf{P})|>1$, then for each $x\in F(\mathbf{P})$, there is a profile $\mathbf{Q}$ of new voters with $|V(\mathbf{Q})|\leq r\cdot|V(\mathbf{P})|$ such that $F(\mathbf{P}+\mathbf{Q})=\{x\}$.
+A natural question is: how small can one make $r$ such that $r$-resolvability is consistent with the other axioms?
+By the proof of Theorem 4 and a scaling argument as in the proof of Theorem 6, $r>1/262$, as $|V(\mathbf{P}_{4}+\mathbf{S}_{2})|\leq 262$ (though this can be improved, e.g., to $1/248$ by Footnote 4).
 Theorem 6.
@@
 Replace each voter in $\mathbf{P}_{1},\dots,\mathbf{P}_{5}$ by $n$ copies of that voter, which results in a multiplication of each margin in $\mathcal{M}_{1},\dots,\mathcal{M}_{5}$ by $n$, and then the proof from Theorem 4 goes through as before but now with $|V(\mathbf{S}_{2})|\leq n$ and $|V(\mathbf{S}_{4})|\leq n$.∎
-These proofs can also be adapted to use negative involvement (Fishburn and Brams 1983) instead of positive involvement: for every profile $\mathbf{P}$, $x\in X(\mathbf{P})$, and profile $\mathbf{P}^{\mathord{\downarrow}x}$ with $V(\mathbf{P})\cap V(\mathbf{P}^{\mathord{\downarrow}x})=\varnothing$ and $X(\mathbf{P})=X(\mathbf{P}^{\mathord{\downarrow}x})$, if $x\not\in F(\mathbf{P})$ and $\mathbf{P}^{\mathord{\downarrow}x}$ consists of a single voter who ranks $x$ uniquely last, then $x\not\in F(\mathbf{P}+\mathbf{P}^{\mathord{\downarrow}x})$.
+These proofs can also be adapted to use negative involvement instead of positive involvement: for every profile $\mathbf{P}$, $x\in X(\mathbf{P})$, and profile $\mathbf{P}^{\mathord{\downarrow}x}$ with $V(\mathbf{P})\cap V(\mathbf{P}^{\mathord{\downarrow}x})=\varnothing$ and $X(\mathbf{P})=X(\mathbf{P}^{\mathord{\downarrow}x})$, if $x\not\in F(\mathbf{P})$ and $\mathbf{P}^{\mathord{\downarrow}x}$ consists of a single voter who ranks $x$ uniquely last, then $x\not\in F(\mathbf{P}+\mathbf{P}^{\mathord{\downarrow}x})$.
+[^8] There was no need for a separate proof for negative involvement in Holliday 2024, since relative to the additional assumption of ordinal margin invariance in that paper, negative involvement is equivalent to positive involvement.
+However, without ordinal margin invariance, they are not equivalent.
+For example, Young’s (1977) method satisfies negative involvement (as well as the Condorcet winner criterion and resolvability) but not positive involvement (see Pérez 2001, Kasper et al.
+2019).
 Contrapositively, if $x\in F(\mathbf{P}+\mathbf{P}^{\mathord{\downarrow}x})$, then $x\in F(\mathbf{P})$.
 Theorem 7.
-There is no voting method satisfying the Condorcet winner and loser criteria, negative involvement, and resolvability.
+For each positive integer $n$, there is no voting method satisfying the Condorcet winner and loser criteria, negative involvement, and $n$-voter resolvability.
 Proof.
-A variant of Lemma 3 holds for negative involvement (cf.
+We give the proof for $n=1$.
+A scaling argument exactly analogous to that in the proof of Theorem 6 handles $n>1$.
+First, observe that a variant of Lemma 3 holds for negative involvement (cf.
 Pérez 2001, Lemma 1): Let $F$ be a voting method satisfying negative involvement and the Condorcet winner criterion.
@@
 Now we adapt the proof of Theorem 4 by adding 147 copies of each linear order of $X(\mathbf{P}_{1})$ to $\mathbf{P}_{1}$, which ensures that the variant of Lemma 3 is applicable when needed in the proof.
-[^6] By starting with at least $147$ copies of each linear order in $\mathbf{P}_{1}$, we ensure that there are at least $121$ copies of each linear order in $\mathbf{P}_{3}+\mathbf{S}_{2}$ and in $\mathbf{P}_{5}+\mathbf{S}_{2}+\mathbf{S}_{4}$, which is greater than the maximal margins in those profiles, so the variant of Lemma 3 is applicable to those profiles.
-Of course, one can use an integer linear program to synthesize a new starting profile with far fewer voters such that all constraints needed for the proof still hold.
+[^9] By starting with at least $147$ copies of each linear order in $\mathbf{P}_{1}$, we ensure that there are at least $121$ copies of each linear order in $\mathbf{P}_{3}+\mathbf{S}_{2}$ and in $\mathbf{P}_{5}+\mathbf{S}_{2}+\mathbf{S}_{4}$, which is greater than the maximal margins in those profiles, so one immediately sees that the variant of Lemma 3 is applicable to those profiles, without any reasoning about the numbers of rankings of specific types (reasoning that could be used to replace $147$ with a smaller number).
+Of course, one can use an integer linear program to synthesize a new starting profile with far fewer voters (indeed, $503$ voters) such that all constraints needed for the proof still hold.
 Then instead of adding 26 voters with the ranking adbecadbec to obtain $\mathbf{P}_{2}$, we remove 26 voters with cebdacebda; instead of removing 7 voters with dacebdaceb from $\mathbf{P}_{2}$ to obtain $\mathbf{P}_{3}$, we add 7 voters with becadbecad; and so on for $\mathbf{P}_{4}$ (removing 23 voters with caedbcaedb) and $\mathbf{P}_{5}$ (adding 7 voters with ecabdecabd).
 The margins of $\mathcal{M}_{1},\dots,\mathcal{M}_{5}$ remain as in the proof of Theorem 4, and then the rest of the proof is the same except that we (i) invoke the variant of Lemma 3 for steps involving defensibility and (ii) use negative involvement instead of positive involvement.∎
-By a scaling argument exactly analogous to that in the proof of Theorem 6, we can extend the proof of Theorem 7 to establish the following.
-Theorem 8.
-For each positive integer $n$, there is no voting method satisfying the Condorcet winner and loser criteria, negative involvement, and $n$-voter resolvability.
-Finally, in response to a previous version of this note with the results above, Dominik Peters (personal communication) suggested that analogous impossibilities might hold with independence of clones (Tideman 1987) instead of the Condorcet loser criterion.
-Indeed, we can confirm that such a swap in Theorems 4 and 6 is possible, as well as in Theorems 78 if we add the axiom of block preservation (Holliday and Pacuit 2025): if the profile $\mathbf{L}$ consists of exactly one copy of each linear order of $X(\mathbf{P})$, then $F(\mathbf{P})\subseteq F(\mathbf{P}+\mathbf{L})$.
-Recall that a set $C\subsetneq X(\mathbf{P})$ with $|C|\geq 2$ is a set of clones in $\mathbf{P}$ if for each $i\in V(\mathbf{P})$ and $x\in X(\mathbf{P})\setminus C$, either $i$ ranks $x$ above each $c\in C$ or $i$ ranks $x$ below each $c\in C$.
+Finally, in response to a previous version of this note with the results above, Dominik Peters (personal communication) suggested that analogous impossibilities might hold with independence of clones (Tideman 1987) in place of the Condorcet loser criterion.
+Indeed, we can confirm that such a swap in Theorems 4 and 6 is possible, as well as in Theorem 7 if we add the axiom of block preservation (Holliday and Pacuit 2025): if the profile $\mathbf{L}$ consists of exactly one copy of each linear order of $X(\mathbf{P})$, then $F(\mathbf{P})\subseteq F(\mathbf{P}+\mathbf{L})$.
+A set $C\subsetneq X(\mathbf{P})$ with $|C|\geq 2$ is a set of clones in $\mathbf{P}$ if for each $i\in V(\mathbf{P})$, $x\in X(\mathbf{P})\setminus C$, and $c,c^{\prime}\in C$, if $i$ ranks $x$ above (resp.
+below) $c$, then $i$ ranks $x$ above (resp.
+below) $c^{\prime}$.
 Given a profile $\mathbf{P}$ and $c\in X(\mathbf{P})$, let $\mathbf{P}_{-c}$ be the profile with $X(\mathbf{P}_{-c})=X(\mathbf{P})\setminus\{c\}$ and $V(\mathbf{P}_{-c})=V(\mathbf{P})$ where for each $i\in V(\mathbf{P})$, $\mathbf{P}_{-c}(i)$ is the restriction of $\mathbf{P}(i)$ to $X(\mathbf{P}_{-c})$.
@@
 Theorem 8.
+  1. 1.
     For each positive integer $n$, there is no voting method satisfying the Condorcet winner criterion, independence of clones, positive involvement, and $n$-voter resolvability.
+  2. 2.
     The same holds with positive involvement replaced by negative involvement plus block preservation.
 Proof.
-We prove the $n=1$ case.
-Reasoning as in the proof of Theorem 6 handles $n>1$.
-For positive involvement, consider a $\mathbf{Q}_{1}$ as in Figure 3, wherein $\{a,b,c,e\}$ is a set of clones.
+We prove the $n=1$ case. Reasoning as in the proof of Theorem 6 handles $n>1$.
+For part 1, consider a $\mathbf{Q}_{1}$ as in Figure 3, wherein $\{a,b,c,e\}$ is a set of clones.
 By Lemma 3, $F(\mathbf{Q}_{1})\subseteq\{a,d\}$.
@@
 The rest of the proof is just like that of Theorem 4 but with $\mathbf{Q}_{i}$ instead of $\mathbf{P}_{i}$.
-For negative involvement, by block preservation, we can add as many blocks of all linear orders as we wish to $\mathbf{Q}_{1}$ to obtain a $\mathbf{Q}_{1}^{+}$ with $F(\mathbf{Q}_{1})\subseteq F(\mathbf{Q}_{1}^{+})\subseteq\{a,d\}$ by the variant of Lemma 3 in the proof of Theorem 7.
+For part 2, by block preservation we can add as many blocks of all linear orders as we wish to $\mathbf{Q}_{1}$ to obtain a $\mathbf{Q}_{1}^{+}$ with $F(\mathbf{Q}_{1})\subseteq F(\mathbf{Q}_{1}^{+})\subseteq\{a,d\}$ by the variant of Lemma 3 in the proof of Theorem 7.
 Hence $F(\mathbf{Q}_{1})\subseteq\{a,d\}$, which implies $a\in F(\mathbf{Q}_{1})$ as above and therefore $a\in F(\mathbf{Q}_{1}^{+})$.
 Then starting from $\mathbf{Q}_{1}^{+}$, we reason with negative involvement instead of positive involvement as in the proof of Theorem 7.∎
+Thus, the tradeoff noted in Section 1 is shown by several inconsistent sets of axioms, summarized in Table 1.
 62 60 42 23 19 3
@@
 Figure 4: margin graphs used in the proof of Theorem 8 with defensible candidates shaded gray.
+Condorcet Condorcet positive negative resolvability independence block
+winner loser involvement involvement of clones preservation
+Theorem 4 $circleddash \circleddash \circleddash$
+Theorem 7 $circleddash \circleddash \circleddash$
+Theorem 8.1 $circleddash \circleddash \circleddash$
+Theorem 8.2 $circleddash \circleddash \circleddash \circleddash$
+Table 1: summary of main impossibility theorems.
+For each theorem, the axioms that are jointly inconsistent are indicated by circled marks in that row.
+In each case, resolvability can be replaced by $n$-voter resolvability.
 Acknowledgements
-I thank Diego Bejarano and Emma Boniface for helpful discussions of the problem posed at the end of Holliday 2024 and Dominik Peters for helpful feedback on this paper.
+I thank Diego Bejarano and Emma Boniface for helpful discussions of the problem posed at the end of Holliday 2024, as well as Dominik Peters and an anonymous referee for helpful feedback on this paper.

### v7

@@
 According to Pérez (2001, p.
-605), these “may be seen as the minimum to require concerning the coherence in the winning set when new voters are added.” [^1] However, a weaker requirement is positive-negative involvement: adding a voter who ranks an initial winner $x$ uniquely first and an initial loser $y$ uniquely last cannot cause $x$ to lose and $y$ to win.
+605), these axioms “may be seen as the minimum to require concerning the coherence in the winning set when new voters are added.” [^1] However, a weaker requirement is positive-negative involvement: adding a voter who ranks an initial winner $x$ uniquely first and an initial loser $y$ uniquely last cannot cause $x$ to lose and $y$ to win.
 Another weaker requirement is bullet-vote positive involvement: adding a voter who ranks an initial winner $x$ uniquely first—and leaves all other candidates unranked or ranks them as tied after xx—cannot cause $x$ to lose.
@@
 Instead of the Condorcet loser criterion, we can use the explicitly variable-candidate axiom of independence of clones (Tideman 1987).
-Finally, for a formalization of the idea that a voting method should not make ties too common, we use the resolvability axiom (Tideman 1986) that if $x$ is initially tied for winning, then $x$ can be made the unique winner by adding a single new voter.
+Finally, for a formalization of the idea that a voting method should not make ties too common, we use the resolvability axiom (Tideman 1986, 1987) that if $x$ is initially tied for winning, then $x$ can be made the unique winner by adding a single new voter.
 As Tideman (1986, p.
@@
   2. 2.
-    $F$ satisfies the Condorcet loser criterion if for every profile $\mathbf{P}$ and $x\in X(\mathbf{P})$, if $\text{Margin}_{\mathbf{P}}(y,x)>0$ for all $y\in X(\mathbf{P})\setminus\{x\}$ (so $x$ is a Condorcet loser), then $x\not\in F(\mathbf{P})$.
+    $F$ satisfies the Condorcet loser criterion if for every profile $\mathbf{P}$ and $x\in X(\mathbf{P})$, if $|X(\mathbf{P})|>1$ and $\text{Margin}_{\mathbf{P}}(y,x)>0$ for all $y\in X(\mathbf{P})\setminus\{x\}$ (so $x$ is a Condorcet loser), then $x\not\in F(\mathbf{P})$.
   3. 3.
@@
 We leave open whether the axioms are jointly satisfiable without ordinal margin invariance if we only ask for a voting method defined on four-candidate profiles.
-If we restrict the domain to two- or three-candidate profiles, then the axioms are jointly satisfied by the Minimax voting method.
-The impossibility proof in the next section makes use of the following concept (cf. Kasper et al. 2019).
+If we restrict the domain to two- or three-candidate profiles, then the axioms are satisfied by the Minimax voting method.
+The impossibility proof in the next section makes use of the following concept (cf. Kasper et al. 2019, Holliday 2024).
 Definition 2.
@@
 Lemma 3 (Moulin 1988, Pérez 1995).
-Let $F$ be a voting method satisfying positive involvement and the Condorcet winner criterion.
+Let $F$ be a voting method satisfying the Condorcet winner criterion and positive involvement.
 Let $\mathbf{P}$ be a profile in which for all $x,x^{\prime},y,y^{\prime}\in X(\mathbf{P})$ with $x\neq x^{\prime}$ and $y\neq y^{\prime}$, if $\text{Margin}_{\mathbf{P}}(x,x^{\prime})>\text{Margin}_{\mathbf{P}}(y,y^{\prime})$, then $\text{Margin}_{\mathbf{P}}(x,x^{\prime})>\text{Margin}_{\mathbf{P}}(y,y^{\prime})+1$.
@@
 After the initial construction of this proof using another version of $\mathcal{M}_{1}$ and different numbers of rankings to add/subtract to obtain an appropriate sequence $\mathcal{M}_{1},\dots,\mathcal{M}_{5}$ of margin graphs for the proof, an integer linear program was used to find a small profile $\mathbf{P}_{1}$ and numbers of rankings to add/subtract to keep the total number of voters down while still satisfying all the constraints needed for each step of the proof.
-Even fewer voters suffice (e.g., $247$ instead of $261$ in $\mathbf{P}_{4}$) if, at the expense of elegance, we allow more distinct ranking types in the profiles.
-Let $\mathbf{P}_{2},\dots,\mathbf{P}_{5}$ be obtained from $\mathbf{P}_{1}$ by adding and subtracting the profiles shown in the equations in Figure 2: $\mathbf{P}_{2}$ is obtained from $\mathbf{P}_{1}$ by adding 26 new voters with the ranking adbecadbec; $\mathbf{P}_{3}$ is obtained from $\mathbf{P}_{2}$ by subtracting $7$ voters with the ranking dacebdaceb, etc.
+Even fewer voters suffice (e.g., $230$ instead of $261$ in $\mathbf{P}_{4}$, the largest $\mathbf{P}_{i}$) if, at the expense of elegance, we allow more distinct ranking types in the profiles.
+Let $\mathbf{P}_{2},\dots,\mathbf{P}_{5}$ be obtained from $\mathbf{P}_{1}$ by adding and subtracting the profiles shown in the equations in Figure 2: $\mathbf{P}_{2}$ is obtained from $\mathbf{P}_{1}$ by adding 26 new voters with the ranking adbecadbec; $\mathbf{P}_{3}$ is obtained from $\mathbf{P}_{2}$ by removing $7$ voters with the ranking dacebdaceb, etc.
 Note there are indeed such $7$ voters in $\mathbf{P}_{1}$, and for the purposes of $\mathbf{P}_{5}$, there are $7$ voters in $\mathbf{P}_{1}$ with the ranking dbacedbace.
@@
 $b$ $d$ $d$ $a$ $e$ $a$
-Figure 1: the profile $\mathbf{P}_{1}$ used in the proof of Theorem 4. The number at the top of each column indicates the number of voters submitting the ranking in that column.
+Figure 1: the profile $\mathbf{P}_{1}$ used in the proof of Theorem 4. The number at the top of each column is the number of voters submitting the ranking in that column.
 $\mathcal{M}_{1}$
@@
 Thus, by the Condorcet loser criterion and Lemma 3, $F(\mathbf{P}_{1})=\{a\}$.
-Since $\mathbf{P}_{2}$ is obtained from $\mathbf{P}_{1}$ by adding voters all of whom rank $a$ uniquely first, $F(\mathbf{P}_{1})=\{a\}$ and repeated application of positive involvement together imply $a\in F(\mathbf{P}_{2})$.
+Since $\mathbf{P}_{2}$ is obtained from $\mathbf{P}_{1}$ by adding voters all of whom rank $a$ uniquely first, together $F(\mathbf{P}_{1})=\{a\}$ and repeated application of positive involvement imply $a\in F(\mathbf{P}_{2})$.
 Then by resolvability, there is some “empty” [^5] Here $\mathbf{S}_{2}$ is not literally an empty profile; if we already have $F(\mathbf{P}_{2})=\{a\}$, then let $\mathbf{P}_{2}+\mathbf{S}_{2}$’ denote $\mathbf{P}_{2}$, $\mathbf{P}_{3}+\mathbf{S}_{2}$’ denote $\mathbf{P}_{3}$, etc.
@@
 or single-voter profile $\mathbf{S}_{2}$ such that $F(\mathbf{P}_{2}+\mathbf{S}_{2})=\{a\}$}.
-Since $\mathbf{P}_{2}+\mathbf{S}_{2}$ is obtained from $\mathbf{P}_{3}+\mathbf{S}_{2}$ by adding voters all of whom rank $d$ uniquely first, $F(\mathbf{P}_{2}+\mathbf{S}_{2})=\{a\}$} and positive involvement together imply $d\not\in F(\mathbf{P}_{3}+\mathbf{S}_{2})$.
+Since $\mathbf{P}_{2}+\mathbf{S}_{2}$ is obtained from $\mathbf{P}_{3}+\mathbf{S}_{2}$ by adding voters all of whom rank $d$ uniquely first, together $F(\mathbf{P}_{2}+\mathbf{S}_{2})=\{a\}$} and positive involvement imply $d\not\in F(\mathbf{P}_{3}+\mathbf{S}_{2})$.
 Since any difference between weights of two distinct edges in $\mathcal{M}_{3}$ is of size at least $4$, no matter which single-voter (or empty) profile $\mathbf{S}_{2}$ is, $b$ is the only defensible candidate besides $d$ in $\mathbf{P}_{3}+\mathbf{S}_{2}$, and $\mathbf{P}_{3}+\mathbf{S}_{2}$ satisfies the margin separation assumption of Lemma 3, so $d\not\in F(\mathbf{P}_{3}+\mathbf{S}_{2})$ implies $F(\mathbf{P}_{3}+\mathbf{S}_{2})=\{b\}$ by Lemma 3.
-Then since $\mathbf{P}_{4}+\mathbf{S}_{2}$ is obtained from $\mathbf{P}_{3}+\mathbf{S}_{2}$ by adding voters all of whom rank $b$ uniquely first, $F(\mathbf{P}_{3}+\mathbf{S}_{2})=\{b\}$ and positive involvement together imply $b\in F(\mathbf{P}_{4}+\mathbf{S}_{2})$.
+Then since $\mathbf{P}_{4}+\mathbf{S}_{2}$ is obtained from $\mathbf{P}_{3}+\mathbf{S}_{2}$ by adding voters all of whom rank $b$ uniquely first, together $F(\mathbf{P}_{3}+\mathbf{S}_{2})=\{b\}$ and positive involvement imply $b\in F(\mathbf{P}_{4}+\mathbf{S}_{2})$.
 Then by resolvability, there is some “empty” or single-voter profile $\mathbf{S}_{4}$ such that $F(\mathbf{P}_{4}+\mathbf{S}_{2}+\mathbf{S}_{4})=\{b\}$.
-Since $\mathbf{P}_{4}+\mathbf{S}_{2}+\mathbf{S}_{4}$ is obtained from $\mathbf{P}_{5}+\mathbf{S}_{2}+\mathbf{S}_{4}$ by adding voters all of whom rank $d$ uniquely first, $F(\mathbf{P}_{4}+\mathbf{S}_{2}+\mathbf{S}_{4})=\{b\}$ and positive involvement together imply $d\not\in F(\mathbf{P}_{5}+\mathbf{S}_{2}+\mathbf{S}_{4})$.
+Since $\mathbf{P}_{4}+\mathbf{S}_{2}+\mathbf{S}_{4}$ is obtained from $\mathbf{P}_{5}+\mathbf{S}_{2}+\mathbf{S}_{4}$ by adding voters all of whom rank $d$ uniquely first, together $F(\mathbf{P}_{4}+\mathbf{S}_{2}+\mathbf{S}_{4})=\{b\}$ and positive involvement imply $d\not\in F(\mathbf{P}_{5}+\mathbf{S}_{2}+\mathbf{S}_{4})$.
 Finally, since any difference between weights of two distinct edges in $\mathcal{M}_{5}$ is of size at least $6$, no matter which single-voter (or empty) profiles $\mathbf{S}_{2}$ and $\mathbf{S}_{4}$ are, $d$ is the only defensible candidate in $\mathbf{P}_{5}+\mathbf{S}_{2}+\mathbf{S}_{4}$, and $\mathbf{P}_{5}+\mathbf{S}_{2}+\mathbf{S}_{4}$ satisfies the margin separation assumption of Lemma 3, so $F(\mathbf{P}_{5}+\mathbf{S}_{2}+\mathbf{S}_{4})=\{d\}$ by Lemma 3.
@@
 Inspection of the proof of Theorem 4 shows that resolvability can be replaced by quasi-resoluteness (in the terminology of Holliday and Pacuit 2023): $|F(\mathbf{P})|=1$ for any profile $\mathbf{P}$ that is uniquely weighted in the sense that for all $x,y,x^{\prime},y^{\prime}\in X(\mathbf{P})$ with $x\neq y$ and $x^{\prime}\neq y^{\prime}$, if $(x,y)\neq(x^{\prime},y^{\prime})$, then $\text{Margin}_{\mathbf{P}}(x,y)\neq\text{Margin}_{\mathbf{P}}(x^{\prime},y^{\prime})$.
-In addition, the package of positive involvement and resolvability can be replaced by strict positive involvement (in the terminology of Dominik Peters, personal communication), which is defined like positive involvement but with $F(\mathbf{P}+\mathbf{P}^{x})=\{x\}$ instead of $x\in F(\mathbf{P}+\mathbf{P}^{x})$.
-[^6] A slight modification of the proof of Theorem 4 shows that positive involvement can be replaced by the weaker axiom that if $F(\mathbf{P})=\{x\}$ and $\mathbf{P}^{x}$ is a profile of new voters all of whom rank $x$ uniquely first, then $x\in F(\mathbf{P}+\mathbf{P}^{x})$ (cf.
-Richelson’s (1978) voter adaptability, which requires that $F(\mathbf{P})=\{x\}$ implies $F(\mathbf{P}+\mathbf{P}^{x})=\{x\}$).
+It also shows that resolvability can be replaced by singleton positive involvement (which Richelson (1978) calls voter adaptability), according to which $F(\mathbf{P})=\{x\}$ implies $F(\mathbf{P}+\mathbf{P}^{x})=\{x\}$ (the analogous point applies to Theorem 7 and what Brandt et al.
+(2025) call singleton negative involvement).
+[^6] A slight modification of the proof of Theorem 4 shows that positive involvement can be replaced by the weaker axiom that if $F(\mathbf{P})=\{x\}$ and $\mathbf{P}^{x}$ is a profile of new voters all of whom rank $x$ uniquely first, then $x\in F(\mathbf{P}+\mathbf{P}^{x})$.
 First, a variant of Lemma 3 holds with the weaker axiom in place of positive involvement, provided we assume $F$ satisfies resolvability and that $\text{Margin}_{\mathbf{P}}(x,x^{\prime})>\text{Margin}_{\mathbf{P}}(y,y^{\prime})$ implies $\text{Margin}_{\mathbf{P}}(x,x^{\prime})>\text{Margin}_{\mathbf{P}}(y,y^{\prime})+3$.
@@
 By the variant of Lemma 3, which is still applicable due to the doubling of margins, we have $F(\mathbf{P}_{5}+\mathbf{S}_{2}+\mathbf{S}_{3}+\mathbf{S}_{4})=\{d\}$.
-Similarly, the negative involvement axiom in Theorem 7 can be replaced by the weaker axiom that if $x\not\in F(\mathbf{P})$ and $\mathbf{P}^{\downarrow x}$ is a profile of new voters all of whom rank $x$ uniquely last, then $F(\mathbf{P}+\mathbf{P}^{\downarrow x})\neq\{x\}$ (cf.
-Brandt et al.’s (2025) singleton negative involvement, which requires that $F(\mathbf{P})\neq\{x\}$ implies $F(\mathbf{P}+\mathbf{P}^{\downarrow x})\neq\{x\}$).
+Similarly, the negative involvement axiom in Theorem 7 can be replaced by the weaker axiom that if $x\not\in F(\mathbf{P})$ and $\mathbf{P}^{\downarrow x}$ is a profile of new voters all of whom rank $x$ uniquely last, then $F(\mathbf{P}+\mathbf{P}^{\downarrow x})\neq\{x\}$.
+This formulation with positive involvement and singleton positive involvement has an implication for probabilistic voting methods if we regard $F(\mathbf{P})$ as the support of a probability distribution: the Condorcet winner and loser criteria are jointly inconsistent with the axiom that adding a voter who ranks $x$ uniquely first never lowers $x$’s probability of winning (from positive to 0 or from $1$ to less than $1$).
 The same proof strategy can be used to show that for each positive integer $n$, resolvability can be replaced by $n$-voter resolvability: for every profile $\mathbf{P}$, if $|F(\mathbf{P})|>1$, then for each $x\in F(\mathbf{P})$, there is a profile $\mathbf{Q}$ with $|V(\mathbf{Q})|\leq n$, $V(\mathbf{P})\cap V(\mathbf{Q})=\varnothing$, and $X(\mathbf{P})=X(\mathbf{Q})$ such that $F(\mathbf{P}+\mathbf{Q})=\{x\}$.
@@
 A natural question is: how small can one make $r$ such that $r$-resolvability is consistent with the other axioms?
-By the proof of Theorem 4 and a scaling argument as in the proof of Theorem 6, $r>1/262$, as $|V(\mathbf{P}_{4}+\mathbf{S}_{2})|\leq 262$ (though this can be improved, e.g., to $1/248$ by Footnote 4).
+By the proof of Theorem 4 and a scaling argument as in the proof of Theorem 6, $r>1/262$, as $|V(\mathbf{P}_{4}+\mathbf{S}_{2})|\leq 262$ (though this can be improved, e.g., to $1/231$ by Footnote 4).
 Theorem 6.
@@
 However, without ordinal margin invariance, they are not equivalent.
-For example, Young’s (1977) method satisfies negative involvement (as well as the Condorcet winner criterion and resolvability) but not positive involvement (see Pérez 2001, Kasper et al.
+For example, Young’s (1977) voting method satisfies negative involvement (as well as the Condorcet winner criterion and resolvability) but not positive involvement (see Pérez 2001, Kasper et al.
 2019).
@@
 [^9] By starting with at least $147$ copies of each linear order in $\mathbf{P}_{1}$, we ensure that there are at least $121$ copies of each linear order in $\mathbf{P}_{3}+\mathbf{S}_{2}$ and in $\mathbf{P}_{5}+\mathbf{S}_{2}+\mathbf{S}_{4}$, which is greater than the maximal margins in those profiles, so one immediately sees that the variant of Lemma 3 is applicable to those profiles, without any reasoning about the numbers of rankings of specific types (reasoning that could be used to replace $147$ with a smaller number).
-Of course, one can use an integer linear program to synthesize a new starting profile with far fewer voters (indeed, $503$ voters) such that all constraints needed for the proof still hold.
+Of course, one can use an integer linear program to synthesize a new starting profile with far fewer voters (indeed, $477$ voters) such that all constraints needed for the proof still hold.
 Then instead of adding 26 voters with the ranking adbecadbec to obtain $\mathbf{P}_{2}$, we remove 26 voters with cebdacebda; instead of removing 7 voters with dacebdaceb from $\mathbf{P}_{2}$ to obtain $\mathbf{P}_{3}$, we add 7 voters with becadbecad; and so on for $\mathbf{P}_{4}$ (removing 23 voters with caedbcaedb) and $\mathbf{P}_{5}$ (adding 7 voters with ecabdecabd).
@@
 Finally, in response to a previous version of this note with the results above, Dominik Peters (personal communication) suggested that analogous impossibilities might hold with independence of clones (Tideman 1987) in place of the Condorcet loser criterion.
-Indeed, we can confirm that such a swap in Theorems 4 and 6 is possible, as well as in Theorem 7 if we add the axiom of block preservation (Holliday and Pacuit 2025): if the profile $\mathbf{L}$ consists of exactly one copy of each linear order of $X(\mathbf{P})$, then $F(\mathbf{P})\subseteq F(\mathbf{P}+\mathbf{L})$.
+Indeed, we will confirm below that such a swap in Theorems 4 and 6 is possible, as well as in Theorem 7 if we add the axiom of block preservation (Holliday and Pacuit 2025): if the profile $\mathbf{L}$ consists of exactly one copy of each linear order of $X(\mathbf{P})$, then $F(\mathbf{P})\subseteq F(\mathbf{P}+\mathbf{L})$.
 A set $C\subsetneq X(\mathbf{P})$ with $|C|\geq 2$ is a set of clones in $\mathbf{P}$ if for each $i\in V(\mathbf{P})$, $x\in X(\mathbf{P})\setminus C$, and $c,c^{\prime}\in C$, if $i$ ranks $x$ above (resp.
@@
 We prove the $n=1$ case. Reasoning as in the proof of Theorem 6 handles $n>1$.
-For part 1, consider a $\mathbf{Q}_{1}$ as in Figure 3, wherein $\{a,b,c,e\}$ is a set of clones.
+For part 1, consider a profile $\mathbf{Q}_{1}$ as in Figure 3, wherein $\{a,b,c,e\}$ is a set of clones.
+Its margin graph is shown as $\mathcal{M}_{1}$ in Figure 4.
 By Lemma 3, $F(\mathbf{Q}_{1})\subseteq\{a,d\}$.
@@
 For part 2, by block preservation we can add as many blocks of all linear orders as we wish to $\mathbf{Q}_{1}$ to obtain a $\mathbf{Q}_{1}^{+}$ with $F(\mathbf{Q}_{1})\subseteq F(\mathbf{Q}_{1}^{+})\subseteq\{a,d\}$ by the variant of Lemma 3 in the proof of Theorem 7.
-Hence $F(\mathbf{Q}_{1})\subseteq\{a,d\}$, which implies $a\in F(\mathbf{Q}_{1})$ as above and therefore $a\in F(\mathbf{Q}_{1}^{+})$.
+Hence $F(\mathbf{Q}_{1})\subseteq\{a,d\}$, which implies $a\in F(\mathbf{Q}_{1})$ as above for part 1 and therefore $a\in F(\mathbf{Q}_{1}^{+})$.
 Then starting from $\mathbf{Q}_{1}^{+}$, we reason with negative involvement instead of positive involvement as in the proof of Theorem 7.∎
@@
 Theorem 8.2 $circleddash \circleddash \circleddash \circleddash$
-Table 1: summary of main impossibility theorems.
-For each theorem, the axioms that are jointly inconsistent are indicated by circled marks in that row.
+Table 1: summary of the main impossibility theorems.
+For each theorem, the axioms that are jointly inconsistent are indicated by circled marks in the row for that theorem.
 In each case, resolvability can be replaced by $n$-voter resolvability.